### v1

# Exploiting Completeness Perception with Diffusion Transformer for Unified 3D MRI Synthesis

Abstract
Missing data problems, such as missing modalities in multi-modal brain MRI and missing slices in cardiac MRI, pose significant challenges in clinical practice. Existing methods rely on external guidance to supply detailed missing state for instructing generative models to synthesize missing MRIs. However, manual indicators are not always available or reliable in real-world scenarios due to the unpredictable nature of clinical environments. Moreover, these explicit masks are not informative enough to provide guidance for improving semantic consistency. In this work, we argue that generative models should infer and recognize missing states in a self-perceptive manner, enabling them to better capture subtle anatomical and pathological variations. Towards this goal, we propose CoPeDiT, a general-purpose latent diffusion model equipped with completeness perception for unified synthesis of 3D MRIs. Specifically, we incorporate dedicated pretext tasks into our tokenizer, CoPeVAE, empowering it to learn completeness-aware discriminative prompts, and design MDiT3D, a specialized diffusion transformer architecture for 3D MRI synthesis, that effectively uses the learned prompts as guidance to enhance semantic consistency in 3D space. Comprehensive evaluations on three large-scale MRI datasets demonstrate that CoPeDiT significantly outperforms state-of-the-art methods, achieving superior robustness, generalizability, and flexibility. The code is available at https://github.com/JK-Liu7/CoPeDiT.
1 Introduction
Magnetic resonance imaging (MRI) is widely used in clinical medicine due to its non-invasive nature and ability to visualize detailed tissue properties [8]. For instance, multi-modal brain MRIs provide complementary information regarding brain anatomy and pathology [27], while volumetric cardiac MRIs reveal detailed cardiac structure and tissue characteristics [7]. However, brain and cardiac MRIs often suffer from missing data in real-world clinical settings [50]. For brain MRIs, certain modalities may be unavailable, while for cardiac MRIs, some slices may be missing, which are commonly caused by limited scanning time, image corruption, and variations in acquisition protocols [33].
Figure 1: The motivation of CoPeDiT. (a) Comparison of completeness perception between CoPeDiT and prior methods. (b) Prompt tokens offer more effective guidance than binary mask codes. (c) CoPeVAE enables more discriminative latent representations. (d) Prompt tokens yield more semantically consistent attention maps in MDiT3D blocks, highlighting stronger interactions between similar modalities (e.g., T1-T1ce, T2-FLAIR).
To address this, generative models have been developed to infer missing data from observed inputs [19, 11]. Existing paradigms rely on auxiliary embeddings, e.g., binary mask codes, as prior knowledge to encode missing patterns (e.g., severity, type, and position) [24, 14, 21, 6, 29]. Nonetheless, this approach has several inherent limitations that hinder its practical applicability. First, in real-world deployments, input images may originate from different hospitals, scanners, or time points. As missing patterns are intrinsically unpredictable, providing manually defined masks in advance is impractical [49, 41]. Second, this over-reliance on predefined masks causes generative models to be unaware of the missing state and unable to distinguish between different modalities or spatial characteristics, thereby reducing robustness to variable and unseen incomplete patterns [22, 20, 51]. Models conditioned on fixed mask representations often suffer from limited generalizability, exhibiting noticeable drop in synthesis quality when encountering unfamiliar missing sites [13, 2, 32]. Finally, binary mask codes lack semantic richness and adaptability, leading to a rigid guidance that weakens spatial alignment and semantic coherence in the synthesized output [37, 17, 44].
Intuitively, generative models should be capable of inferring and detecting the incomplete state spontaneously, rather than relying on externally manual guidance [17, 12]. Motivated by this, we pose a central question: ‘Can we empower the model with the ability to perceive missing states on its own?’ In light of this, we exploit a long neglected and underexplored property of generative models in medical imaging, i.e., completeness perception, to enhance flexibility and generalizability under arbitrary missing MRI conditions, as illustrated in Fig. 1 (a). Our fundamental insight is to enable the generative model to recognize the fine-grained incomplete state information in a self-perceptive manner, and to leverage this understanding as internal prompts to guide the generation process. We hypothesize that, for diffusion models, such self-guided prompts may serve as an effective alternative to manually defined masks, and potentially offer even stronger guidance signals (Fig. 1 (b)). The main reason lies in the fact that this self-perceptive strategy encourages the model to learn both global and local anatomical structures and lesion patterns at coarse and fine levels, thus enabling more semantically coherent generation of the missing MRI regions during synthesis (Fig. 1 (c)(d)) [23, 18].
Driven by our motivation, we propose CoPeDiT, a 3D latent diffusion model (LDM) framework for unified 3D MRI synthesis. Technically, our framework builds on two core components: (i) Unlike prior approaches that require explicit missing indicators, a novel tokenizer with a completeness perception function, CoPeVAE, is proposed to autonomously assess the integrity of modalities or volumes through tailored self-supervised pretext tasks. By detecting anatomical structures and variations in lesion patterns, CoPeVAE develops a comprehensive understanding of 3D MRIs. This enables CoPeDiT to eliminate the need for manual intervention with flexible adaptability, enhancing the method’s autonomy and improves its feasibility for real-world clinical deployment with diverse missing patterns. (ii) A novel diffusion transformer architecture, MDiT3D, is developed to unlock the potential of Diffusion Transformers (DiT) [34] for medical image synthesis, offering a new paradigm for generating high-quality volumetric data. MDiT3D leverages tokenization and attention mechanisms to model long-range and irregular contextual dependencies in high-dimensional, structurally complex 3D MRIs. It supports conditional generation guided by completeness-aware prompts, which provide semantic-meaningful guidance on the synthesis process. The learned prompt tokens comprehensively absorb information from available modalities or slices during the diffusion process and guide the generation of the missing ones. Incorporating the above two innovations, our architecture not only enables adaptive self-guidance synthesis in a mask-free manner which exhibits strong robustness and generalizability, but also demonstrates improved structural coherence and enhanced preservation of fine-grained anatomical details.
Our main contributions are summarized as follows:
  • We propose a unified framework, dubbed CoPeDiT, for both 3D brain and cardiac missing MRI synthesis under arbitrary incomplete scenarios, without the need for explicit external indicators as guidance.
  • We empower our tokenizer, CoPeVAE, with a strong capacity to perceive completeness by seamlessly integrating carefully designed pretext tasks, enabling the model to recognize missing states and learn informative, self-guided prompts.
  • We present MDiT3D, a novel latent diffusion transformer for 3D MRI synthesis, featuring a dedicated architecture for capturing modality/spatial relationships and leveraging generated prompts as conditional guidance.
  • Extensive experiments on three datasets demonstrate that our model surpasses state-of-the-art (SOTA) methods, achieving generalizability, robustness, and real-world clinical applicability.
2 Related Work
Figure 2: The overview framework of CoPeVAE. We implement two variants, CoPeVAE-B and CoPeVAE-C, with slight architectural modifications for brain and cardiac MRI synthesis tasks, respectively.
MRI Synthesis. Recent efforts focus on unified MRI synthesis using generative models such as Generative Adversarial Networks (GAN) and diffusion models [52, 56, 1, 36]. For example, MMGAN [43] employs a single network to handle diverse missing configurations, while Hyper-GAE [53] introduces a shared hyper-encoder and decoder architecture with a graph-based fusion module for reconstructing missing brain MRIs. MMT [24] proposes to leverage Swin Transformers [25] to capture inter-modal dependencies. M2DN [29] presents a diffusion-based method that jointly models all modalities using a multi-input multi-output framework. Despite these advances, most existing methods still rely on externally provided masks to encode missing patterns in randomly incomplete cases. In contrast, CoPeDiT explores self-perceptive capability of generative models to autonomously recognize data completeness, enabling more flexible and high-fidelity MRI synthesis.
Diffusion Models. Diffusion models have shown strong generative capacities across various computer vision tasks [26, 28, 55]. Denoising Diffusion Probabilistic Models [16] generate images by gradually adding Gaussian noise and learning to reverse the process. LDM [39] improve efficiency by operating in a compressed latent space using autoencoders. Recently, Diffusion transformers [34] have been effective alternatives to U-Net [40] in LDMs, achieving competitive performance in natural image generation. However, most prior medical image generation approaches still rely on U-Net-based architectures as the diffusion backbone, leaving the potential of transformer-based diffusion models underexplored in this domain [48, 31]. In this paper, we devise MDiT3D, by replacing the U-Net backbone with a diffusion transformer and incorporating task-specific architectural modifications, MDiT3D is capable of generalizing to 3D MRI data and capturing long-range dependencies of high-dimensional 3D MRIs.
Figure 3: The overview framework of MDiT3D. PE and RoPE denote positional embeddings and rotary position embeddings [45, 54], respectively.
3 Methodology
3.1 Notations
We consider two 3D MRI synthesis tasks: brain missing modality synthesis and cardiac missing slice synthesis. To unify the formulation, let $\mathcal{M}=\{\mathbf{x}^{i}\}_{i=1}^{m}$ denote a complete MRI sample, where $m$ is the total number of modalities or slices. The available and missing subsets are denoted as $\mathcal{M}^{S}=\{\mathbf{x}^{s_{i}}\}_{i=1}^{s}$ and $\mathcal{M}^{C}=\{\mathbf{x}^{c_{i}}\}_{i=1}^{c}$, respectively, with $m=s+c$. (1) Brain missing modality synthesis: $\mathcal{M}$ represents all brain MRIs with $m$ modalities, and the task is to infer $\mathcal{M}^{C}$ from $\mathcal{M}^{S}$. (2) Cardiac missing slice synthesis: $\mathcal{M}$ corresponds to a volume with $m$ slices, where missing slices $\mathcal{M}^{C}$ are consecutive, non-overlapping with $\mathcal{M}^{S}$, and are to be reconstructed from the available ones. Notably, our experimental setup considers randomly generated incomplete cases from all the possible combinations as input, with a varying number and length of missing modalities or slices, aiming to mimic real-world clinical environments.
3.2 Stage I: Completeness Perception Tokenizer
The core idea of CoPeVAE (Fig. 2) is that detecting the completeness of high-resolution MRI data enforces the model to perceive both global anatomy and local lesion patterns, thereby producing high-quality prompts as diffusion guidance. Building upon VQGAN [47, 10], we deploy a 3D autoencoder jointly trained with self-supervised pretext tasks. Each task employs a prompt encoder, denoted as $\mathcal{F}_{1},\mathcal{F}_{2}$ and $\mathcal{F}_{3}$, to transform latent tokens (learned by the encoder $\mathcal{E}$) into low-dimensional prompt tokens. All prompt encoders contain 3D Conv layers followed by spatial average pooling. Afterwards, task-specific projection heads are applied for multi-granular classification and contrastive learning.
Data Sampling. To improve CoPeVAE’s adaptability to diverse missing cases, we employ a dual-random sampling strategy, where both the number/length of missing elements and the modality types/slice positions are randomly sampled. Given a complete set $\mathcal{M}$, we first randomly sample a missing count $c\in\{1,\dots,m-1\}$, then uniformly select $c$ elements to form the missing subset $\mathcal{M}^{C}$, with the remainder forming the incomplete subset $\mathcal{M}^{S}$.
Task 1. Missing Number/Length Detection. Equipped with global contextual awareness, the tokenizer is capable of answering the question that how many modalities/slices are missing in the incomplete input. This task aims to enable the tokenizer to identify the severity of incompleteness by perceiving the global context of MRIs, thereby learning modality/spatial attributes in a coarse-grained manner. We formulate this task as an (m1)(m-1)-class classification task and define the loss using the cross-entropy as
$\small\mathcal{L}_{d}=\mathcal{L}_{\text{cls}}(\mathcal{H}_{1}(\mathcal{F}_{1}(\mathbf{z}^{s}))_{c}).$ (1)
where $\mathcal{F}_{1}$ and $\mathcal{H}_{1}$ represent the prompt encoder and the projection head for missing detection, respectively. The learned prompt tokens $\mathbf{p}^{d}=\mathcal{F}_{1}(\mathbf{z}^{s})$ are rich in information about the severity of missing state.
Task 2. Incompleteness Positioning. By identifying which modalities or slices are missing, CoPeVAE yields prompt tokens $\mathbf{p}^{p}=\mathcal{F}_{2}(\mathbf{z}^{s})$ that capture semantically meaningful local properties. The motivation of this task is to drive the model to develop a finer-grained contextual understanding of subtle anatomical structures and detailed pattern variations. Although the missing position implicitly encodes the count, Task 1 learns a modality/slice-agnostic global magnitude prior that calibrates the conditioning strength, while this task provides discrete, spatially localized cues about the exact missing identity. The incorporation of the two tasks improves robustness to errors from either signal. This task is formulated as an $m$-class classification problem and is also optimized using the cross-entropy loss, defined as follows:
$\small\mathcal{L}_{p}=\mathcal{L}_{\text{cls}}(\mathcal{H}_{2}(\mathcal{F}_{2}(\mathbf{z}^{s}))_{I}).$ (2)
where $I$ denotes the index of missing type/position.
Task 3. Missing Modality/Slice Assessment. Motivated by the observation that modalities or slices from the same scan share more similar anatomical and textural context than those from different scans, we adopt an inter-modal/inter-slice contrastive learning scheme [38] to serve as a missing data estimator. Specifically, we take the incomplete latent tokens $\mathbf{z}^{s}$ as anchor, the corresponding missing latent $\mathbf{z}^{c}$ from the same subject as positives, and latent tokens $\mathbf{z}^{c}_{-}$ from different subjects as negatives. The contrastive loss is defined as follows:
$\mathcal{L}_{c}=-\log\frac{\varphi\left(\mathcal{H}_{3}(\mathbf{p}^{s}),\mathcal{H}_{3}(\mathbf{p}^{c})\right)}{\varphi\left(\mathcal{H}_{3}(\mathbf{p}^{s}),\mathcal{H}_{3}(\mathbf{p}^{c})\right)+\sum\varphi\left(\mathcal{H}_{3}(\mathbf{p}^{s}),\mathcal{H}_{3}(\mathbf{p}^{c}_{-})\right)}.$
where $\mathbf{p}^{s}=\mathcal{F}_{3}(\mathbf{z}^{s})$, $\quad\mathbf{p}^{c}=\mathcal{F}_{3}(\mathbf{z}^{c})$, $\varphi(a,b)=\exp(a\cdot b/\tau)$, and $\tau$ is the temperature parameter. This contrastive learning scheme encourages the model to focus on inter-modal/slice contextual differences, thus improving anatomical coherence and fine-grained detail preservation. The overall loss of CoPeVAE is formulated as
$\small\mathcal{L}_{\text{tok}}=\mathcal{L}_{\text{rec}}+\lambda(\mathcal{L}_{d}+\mathcal{L}_{p}+\mathcal{L}_{c}).$ (4)
where $\lambda$ is the weighting coefficient, and $\mathcal{L}_{\text{rec}}$ denotes the reconstruction loss that consists of a pixel-wise L1 loss, vector quantization loss, adversarial loss, and perceptual loss.
Table 1: Quantitative results for multi-modal brain MRI synthesis on BraTS and IXI datasets. The numbers of the second row denote the number of missing modalities. The best performances are bolded.
BraTS IXI
1 2 3 1 2
[…]
GAN-based Methods
MMGAN [43] 24.71 $\pm$1.57 0.817 $\pm$0.027 27.94 $\pm$1.05 24.38 $\pm$1.74 0.806 $\pm$0.023 32.48 $\pm$1.66 24.06 $\pm$1.92 0.794 $\pm$0.031 39.37 $\pm$2.10 22.29 $\pm$1.35 0.684 $\pm$0.015 70.91 $\pm$3.64 21.13 $\pm$1.49 0.668 $\pm$0.019 93.57 $\pm$3.42
MMT [24] 25.19 $\pm$1.41 0.824 $\pm$0.017 24.53 $\pm$1.33 24.50 $\pm$1.55 0.811 $\pm$0.020 29.57 $\pm$1.82 23.92 $\pm$1.53 0.801 $\pm$0.021 39.66 $\pm$1.95 22.64 $\pm$1.49 0.698 $\pm$0.013 53.60 $\pm$3.14 21.82 $\pm$1.60 0.687 $\pm$0.019 72.24 $\pm$2.98
Hyper-GAE [53] 24.65 $\pm$1.62 0.813 $\pm$0.022 28.97 $\pm$1.57 24.42 $\pm$1.74 0.808 $\pm$0.024 33.52 $\pm$1.78 23.86 $\pm$1.88 0.788 $\pm$0.029 41.79 $\pm$2.21 22.12 $\pm$1.33 0.682 $\pm$0.017 72.62 $\pm$3.78 20.91 $\pm$1.45 0.662 $\pm$0.021 98.79 $\pm$4.05
Diffusion Model-based Methods
LDM [39] 23.84 $\pm$1.52 0.805 $\pm$0.019 36.47 $\pm$1.54 23.12 $\pm$1.64 0.798 $\pm$0.021 45.93 $\pm$2.41 22.65 $\pm$1.77 0.791 $\pm$0.025 52.50 $\pm$2.07 21.36 $\pm$1.28 0.679 $\pm$0.016 86.62 $\pm$3.86 20.94 $\pm$1.53 0.654 $\pm$0.020 112.57 $\pm$5.02
ControlNet [58] 23.98 $\pm$1.49 0.808 $\pm$0.018 34.09 $\pm$1.93 23.34 $\pm$1.60 0.801 $\pm$0.020 41.28 $\pm$2.49 22.85 $\pm$1.68 0.795 $\pm$0.022 48.07 $\pm$1.99 21.83 $\pm$1.51 0.681 $\pm$0.015 81.26 $\pm$3.85 21.07 $\pm$1.50 0.661 $\pm$0.019 103.77 $\pm$5.17
M2DN [29] 26.45 $\pm$1.36 0.830 $\pm$0.016 21.29 $\pm$1.18 25.87 $\pm$1.48 0.820 $\pm$0.017 27.36 $\pm$1.62 25.21 $\pm$1.59 0.809 $\pm$0.024 32.40 $\pm$1.93 23.47 $\pm$1.43 0.715 $\pm$0.014 42.52 $\pm$3.08 22.81 $\pm$1.56 0.702 $\pm$0.018 55.64 $\pm$3.62
CoPeDiT 28.26 $\pm$1.24 0.842 $\pm$0.019 12.67 $\pm$0.98 28.13 $\pm$1.49 0.831 $\pm$0.021 13.25 $\pm$1.45 27.91 $\pm$1.41 0.822 $\pm$0.023 14.89 $\pm$1.61 24.34 $\pm$1.21 0.732 $\pm$0.016 25.84 $\pm$2.58 23.92 $\pm$1.53 0.721 $\pm$0.020 32.53 $\pm$3.16
3.3 Stage II: 3D MRI Diffusion Transformer
We propose MDiT3D (see Fig. 3), extending the model configuration of DiT [34] with several key design innovations to enable MDiT3D to adapt to 3D modalities and volumetric MRI. It takes the high-dimensional latent tokens $\mathbf{z}^{s}$ as input, along with the learned prompt tokens $\mathbf{p}=\mathbf{p}^{d}\,\|\,\mathbf{p}^{p}\,\|\,\mathbf{p}^{s}$ from our CoPeVAE tokenizer, which encode comprehensive missing state information and serve as conditional guidance for synthesizing. Without loss of generality, we use MDiT3D-B as a representative configuration to illustrate the architectural design.
3D Patchify & 3D Positional Embeddings. We apply a 3D patchify operator with patch size $p\times p\times p$ to obtain $\mathbf{h}\in\mathbb{R}^{m\times T\times d}$ [30], where $m$ denotes the number of modality, and $d$ is the embedding dimension. Unlike the standard 2D positional embeddings [9] used in DiT, we leverage 3D frequency-based sine-cosine positional embeddings (3D PE) to better encode spatial positional information within the 3D MRI data.
Alternating Blocks & Prompt Injection. Two distinct types of blocks are designed tailored for brain and cardiac tasks, respectively. For the brain MRI synthesis task, we adopt spatial and modal blocks: the former focuses on capturing 3D spatial information, while the latter models inter-modal relationships. Before each modal block, the tokens $\mathbf{h}$ are reshaped into $\mathbb{R}^{T\times m\times d}$ to facilitate inter-modal interaction, then reshaped back for subsequent spatial processing. These two types of blocks are stacked in an alternating fashion to form a total of $N$ blocks. The motivation behind this design lies in the need for the model to capture both inter- and intra-modal relationships to boost contextual understanding and anatomical consistency. Additionally, adaptive layer normalization is employed to inject prompts into the blocks. Notably, inspired by the intuition that conditioning signals should be subtle yet informative, prompt tokens are injected only into the modal blocks.
Joint Reconstruction & Synthesis. We concatenate the available and the noisy latent (corresponding to missing sections) as the input to MDiT3D. Moreover, a joint reconstruction & synthesis scheme is introduced, where only the missing sections undergo noise injection and denoising during the diffusion process, while the observed regions remain unchanged, serving as auxiliary context that provides rich semantic guidance for generation. Given that directly synthesizing the entire modality/volume is more informative and efficient than simply predicting zero noise for the available latent regions [29], we adopt x-prediction loss as objective function for MDiT3D, as below:
$\small\mathcal{L}_{\text{diff}}=\mathbb{E}_{\mathbf{z}_{0},\epsilon,t}\left[\left\|\mathbf{z}_{0}-f_{\theta}(\mathbf{z}_{t},t,\mathbf{p})\right\|_{2}^{2}\right].$ (5)
where $\mathbf{z}_{0}=\mathbf{z}^{s}\,\|\,\mathbf{z}^{c}$ denotes the concatenated clean latent tokens, $\mathbf{z}_{t}=\mathbf{z}^{s}\,\|\,\mathbf{z}_{t}^{c}$ is the noisy input at timestep $t$, $\mathbf{p}$ is the prompt tokens, and $f_{\theta}$ represents our MDiT3D model.
4 Experiments and Results
Table 2: Quantitative results for Cardiac MRI synthesis on UKBB dataset. The numbers of the second row denote the length of missing slices. The best performances are bolded.
UKBB
8 16 24
[…]
GAN-based Methods
MMGAN [43] 25.81 $\pm$0.86 0.815 $\pm$0.012 16.68 $\pm$0.93 24.35 $\pm$0.84 0.793 $\pm$0.009 27.52 $\pm$0.90 23.06 $\pm$1.08 0.776 $\pm$0.018 45.88 $\pm$1.32
MMT [24] 26.02 $\pm$0.82 0.824 $\pm$0.009 15.90 $\pm$0.88 24.73 $\pm$0.87 0.809 $\pm$0.010 23.75 $\pm$0.97 24.12 $\pm$1.06 0.794 $\pm$0.017 37.39 $\pm$1.27
Hyper-GAE [53] 25.23 $\pm$0.89 0.810 $\pm$0.012 19.02 $\pm$0.96 23.66 $\pm$0.94 0.789 $\pm$0.013 31.84 $\pm$1.15 22.70 $\pm$1.14 0.771 $\pm$0.019 48.27 $\pm$1.39
Diffusion Model-based Methods
LDM [39] 24.17 $\pm$0.96 0.795 $\pm$0.012 24.61 $\pm$0.94 23.04 $\pm$1.03 0.778 $\pm$0.010 44.93 $\pm$1.49 22.19 $\pm$1.21 0.761 $\pm$0.018 60.02 $\pm$1.68
ControlNet [58] 24.62 $\pm$0.91 0.801 $\pm$0.012 22.47 $\pm$1.06 23.30 $\pm$0.98 0.784 $\pm$0.014 37.51 $\pm$1.55 22.46 $\pm$1.16 0.765 $\pm$0.020 54.93 $\pm$1.74
M2DN [29] 25.48 $\pm$0.79 0.814 $\pm$0.011 17.72 $\pm$0.85 24.62 $\pm$0.87 0.803 $\pm$0.011 24.15 $\pm$1.06 24.03 $\pm$0.99 0.780 $\pm$0.021 40.08 $\pm$1.59
CoPeDiT 26.42 $\pm$0.81 0.831 $\pm$0.013 15.53 $\pm$0.88 26.07 $\pm$0.74 0.826 $\pm$0.013 18.21 $\pm$0.90 25.39 $\pm$0.86 0.817 $\pm$0.016 25.84 $\pm$1.22
Figure 4: Qualitative results on BraTS dataset. The results are depicted in the axial (top), sagittal (middle) and coronal (bottom) views of the 3D MRI volume. The visual results on IXI dataset are provided in the supplementary material.
Figure 5: Qualitative results of comparison on UKBB dataset. The top and bottom results correspond to the first and last missing slices within a given volume, respectively.
Table 3: Ablation study on the contribution of each pretext task to CoPeVAE’s reconstruction capacity. ‘Cardiac’ refers to the aggregated cardiac MRI dataset consisting of four sources, which are used for training our CoPeVAE model.
BraTS IXI Cardiac
[…]
w/o Task 1 34.38 0.926 30.92 0.914 32.34 0.921
w/o Task 2 33.62 0.918 30.35 0.908 31.59 0.914
w/o Task 3 34.69 0.929 31.13 0.917 32.62 0.925
CoPeVAE 35.05 0.935 31.28 0.921 33.34 0.931
Table 4: Ablation study on the contribution of completeness-aware prompt tokens.
BraTS UKBB
1 2 3 8 24
[…]
w/o $\mathbf{p}^{d}$ 27.35 0.833 16.04 27.02 0.824 17.30 26.73 0.814 19.62 25.82 0.819 16.34 24.53 0.803 33.80
w/o $\mathbf{p}^{p}$ 26.92 0.829 18.46 26.43 0.819 20.23 26.06 0.810 25.13 25.17 0.812 20.08 24.39 0.798 36.54
w/o $\mathbf{p}^{s}$ 27.56 0.835 15.26 27.22 0.827 16.37 26.90 0.815 17.58 26.27 0.828 16.79 25.08 0.813 29.83
w/o Prompts 25.92 0.823 25.69 25.06 0.807 32.17 24.83 0.802 37.97 24.70 0.797 23.29 23.56 0.778 42.17
w/ Mask Codes 27.18 0.831 17.42 26.82 0.816 20.06 26.18 0.809 24.59 26.15 0.823 18.15 24.65 0.802 35.86
CoPeDiT 28.26 0.842 12.67 28.13 0.831 13.25 27.91 0.822 14.89 26.42 0.831 15.53 25.39 0.817 25.84
Table 5: Quantitative results by incorporating our learned prompt tokens into baselines instead of mask codes.
BraTS
[…]
MMT [24] 25.19 0.824 24.53
+ Prompts (ours) 25.68 (+0.49) 0.826 (+0.002) 22.07 (–2.46)
Hyper-GAE [53] 24.65 0.813 28.97
+ Prompts (ours) 25.26 (+0.61) 0.822 (+0.009) 24.26 (–4.71)
M2DN [29] 26.45 0.830 21.29
+ Prompts (ours) 27.23 (+0.78) 0.837 (+0.007) 17.06 (–4.23)
4.1 Experimental Setup
Datasets. (i) Brain MRI Datasets. We evaluate the effectiveness of CoPeDiT on two public brain MRI datasets: BraTS 2021 [3] and IXI [5]. The BraTS 2021 dataset includes 1251 subjects with multi-modal MRI scans across four modalities: T1, T1ce, T2, and FLAIR. The IXI dataset contains 577 subjects with three MRI modalities: T1, T2, and PD. (ii) Cardiac MRI Datasets. Missing slice synthesis experiments are conducted on four cardiac MRI datasets: UK Biobank (UKBB) [35], MESA [57], ACDC [4], and MSCMR [59]. The model is trained on the combined dataset including all four sources with 32,248 MRI volumes in total, while performance comparisons are conducted on the UKBB dataset. We randomly select 80% of the data for training and use the remaining 20% as the test set. Please refer to the supplementary material for more dataset details.
Implementation Details. The compression rate of CoPeVAE is set to $(8,8,8)$. For hyperparameters, the dimension of each prompt token is set to 512, $\tau$ and $\lambda$ are set to 0.2 and 1e-2, respectively. The model is trained with a global batch size of 8 for CoPeVAE-B and 64 for CoPeVAE-C, using a learning rate of 1e-4. Regarding MDiT3D, we set the number of time steps to 500 with linearly scaled noise scheduling. The model is trained for 100k steps with a global batch size of 32 and learning rate of 5e-5. All training is conducted on four NVIDIA A100 GPUs. More details are provided in the supplementary material.
4.2 Performance Comparison
Quantitative Results. We compare our method against six recent SOTA models for 3D brain and cardiac MRI synthesis, including three GAN-based approaches [43, 24, 53] and three advanced diffusion model-based methods [39, 58, 29]. We report results by reimplementing all baselines from scratch under identical experimental settings to ensure fair comparisons. Performance is evaluated using three metrics: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM), and Fréchet Inception Distance (FID) [15] that uses Inception-V3 [46] as the feature extractor. Quantitative results on the three datasets are presented in Tables 1 and  2, where the second row indicates the number of missing modalities or slices $c$, respectively. The results show that CoPeDiT consistently outperforms all baselines across all missing configurations in both synthesis tasks, showcasing superior generalizability. Notably, the performance gains are more pronounced in scenarios with a larger number of missing modalities/slices, highlighting the robustness of our completeness-aware prompts in handling complex missing cases. Besides, CoPeDiT achieves significantly lower FID, indicating that the generated MRIs are more anatomically coherent and texture-preserving, with feature distributions that are closer to real clinical data, thus enhancing perceptual realism and diagnostic plausibility.
Qualitative Results. As depicted in Fig. 4 and 5, our model generates synthetic MRIs that exhibit the highest visual similarity to the ground truth images, particularly in accurately capturing tumor regions. Our CoPeDiT excels at preserving subtle textural details and modeling complex anatomical structures within brain tissues, justifying our motivation that incorporating completeness perception leads to improved anatomical consistency and realism.
Table 6: Ablation study on the choice of prompt injection positions within the MDiT3D blocks. ‘Spatial / Planar’ indicates that prompt tokens are injected into the spatial and planar blocks for brain and cardiac tasks, respectively. ‘Both’ indicates that prompt tokens are injected into both the spatial and modal blocks for brain tasks, and into both the planar and spatial blocks for cardiac tasks. ‘Modal / Spatial’ indicates that prompt tokens are injected into the modal and spatial blocks, respectively.
BraTS UKBB
1 2 3 8 24
[…]
Spatial / Planar 27.47 0.834 15.41 27.20 0.825 16.83 27.03 0.817 17.45 26.29 0.827 16.82 25.22 0.814 29.24
Both 27.76 0.836 13.94 27.54 0.827 15.48 27.29 0.818 16.77 26.33 0.827 16.39 25.28 0.815 28.45
Modal / Spatial 28.26 0.842 12.67 28.13 0.831 13.25 27.91 0.822 14.89 26.42 0.831 15.53 25.39 0.817 25.84
Table 7: Ablation study on the contribution of 3D MRI Diffusion Transformer.
BraTS UKBB
1 2 3 8 24
[…]
UNet [39] 25.61 0.825 23.56 25.20 0.816 28.37 24.78 0.805 36.70 24.46 0.800 22.34 23.01 0.778 41.59
DiT [34] 26.78 0.835 19.72 26.14 0.823 21.84 25.74 0.808 24.47 25.89 0.825 16.28 24.78 0.799 36.10
DiT-3D [30] 26.84 0.835 19.23 26.23 0.822 21.95 25.90 0.810 26.08 26.08 0.825 16.37 24.83 0.806 33.73
MDiT3D 28.26 0.842 12.67 28.13 0.831 13.25 27.91 0.822 14.89 26.42 0.831 15.53 25.39 0.817 25.84
4.3 Ablation Study
Effect of Pretext Tasks. As Table 3 displays, CoPeVAE preserves strong reconstruction capability while benefiting from the incorporation of pretext tasks. Each task contributes positively, and their combination leads to further improvements. This improvement can be attributed to the fact that pretext tasks promote the tokenizer to capture anatomical structure variances in both coarse and fine-grained manner, learning highly discriminative features.
Effect of Prompt Tokens. Table 4 presents the synthesis performance associated with each learned prompt token and mask code. The mask codes are generated as one-hot vectors whose dimension corresponds to the number of modalities or slices, indicating the corresponding missing positions. Notably, we employ an identical injection strategy for all prompts and masks to ensure a fair comparison. As shown, all individual prompt tokens and their combinations improve model performance, surpassing guidance based on binary mask codes. These findings strongly highlight the effectiveness and rationale of our completeness perception scheme. The incompleteness positioning task yields the most effective prompts, likely due to its explicit emphasis on positioning missing sections along the modality/slice dimension, thus increasing the sensitivity to subtle structural variations. To further explore the potential of our learned prompts, we incorporate them into prior baselines as replacements for mask codes. As illustrated in Table 13, this substitution consistently improves their performance, demonstrating the utility and generalizability of our completeness-aware prompts, which can serve as a flexible paradigm for improving emerging MRI synthesis methods.
Choice of Prompt Injection Position. This experiment aims to justify our choice of injecting prompts exclusively into the modal and spatial blocks for brain and cardiac tasks, respectively. As illustrated in Table 14, injecting prompts into the modal/spatial block yields the best overall performance, whereas injecting into spatial/planar blocks or both blocks results in inferior outcomes. This can be attributed to the fact that aligning the prompt with the block that explicitly models the task’s primary dependency, namely modality fusion for brain and through-plane continuity for cardiac, maximizes the efficacy of conditional signals.
Impact of 3D MRI Diffusion Transformer. The quantitative evaluation of MDiT3D and existing diffusion networks [39, 34, 30] is presented in Table 7. The results illustrate that MDiT3D consistently achieves superior performance, confirming that the specialized architectural designs effectively enhance the quality of 3D MRI synthesis.
4.4 Visualization Analysis
Figure 6: Visualization of salient regions on BraTS dataset.
Figure 7: t-SNE visualization on the BraTS dataset of (a) The count-focused prompt $\mathbf{p}^{d}$ colored by missing-number classes; and (b) The identity-focused prompt $\mathbf{p}^{p}$ colored by missing modality.
Figure 8: Visualization of attention maps from individual modal blocks on BraTS dataset under a missing configuration of [1, 0, 0, 0], where the first modality is absent.
Salient Regions. To better understand the learning procedure of pretext tasks, we visualize their activation maps using GradCAM [42]. As shown in Fig. 6, the four examples reveal strong correlations between salient regions and modality-discriminative features. Specifically, these patterns mirror each task’s objective: Task 1 and Task 3 must assess global consistency, so they rely on coarse anatomical layout (gray and white matter) that summarize the volume. In contrast, Task 2 needs to pinpoint where/which is missing, so it keys on high-frequency, modality-specific cues (tumors, lesions, and white matter hyperintensities in FLAIR) to localize. The results emphasize that our pretext tasks effectively capture modality-specific properties, understanding the MRI context and relationships.
Prompt Distribution. We implement the t-SNE visualization of learned prompts on BraTS and color them by ground-truth incompleteness labels. As depicted in Fig. 7, the count-focused prompt $\mathbf{p}^{d}$ forms compact, well-separated clusters aligned with the missing number classes corresponding to each missing state (1/2/3 absent modalities), while the identity-focused prompt $\mathbf{p}^{p}$ produces modality-specific clusters (T1, T1ce, T2, FLAIR) with clear boundaries, highlighting that the two prompts capture complementary aspects of incompleteness.
Attention Maps. We investigate the contribution of prompts by visualizing the attention maps of each modal block. As shown in Fig. 8, our prompts guide the attention mechanism to progressively focus on the actual missing elements (the first row and column) with the increase of block layers. In comparison, mask codes lack sufficient informativeness and fail to provide nuanced guidance that aligns with the true missing state.
5 Conclusion
This work presents CoPeDiT, a unified model for 3D brain and cardiac MRI synthesis that explores completeness perception. We demonstrate that enabling the model to autonomously infer missing state, rather than relying on externally pre-defined masks, can provide more discriminative and informative guidance. To this end, we equip our tokenizer with completeness perception capability through elaborated pretext tasks. MDiT3D is then developed to utilize the learned prompt tokens as guidance for 3D MRI generation. Experimental evaluation validates the superiority of CoPeDiT in terms of accuracy, generalizability, robustness, and adaptability across diverse missing scenarios. Furthermore, CoPeDiT enables mask-free synthesis through perceiving completeness state, enhancing flexibility and applicability in practical clinical settings.
Supplementary Material
Appendix A Datasets
The details of the brain and cardiac MRI datasets used in our experiment are summarized in Table 8. Notably, we train our CoPeVAE and MDiT3D models on the brain MRI synthesis task on BraTS and IXI datasets separately, due to differences in the number and types of modalities. The evaluation and results are also reported for the two datasets separately. For Cardiac MRI synthesis, we leverage a combination of all four datasets to train both two-stage models.
Table 8: Details of brain and cardiac MRI datasets.
Datasets Modality Cases Train Test
Brain MRI
BraTS [3] T1, T1ce, T2, FLAIR 1251 1000 251
IXI [5] T1, T2, PD 577 462 115
Cardiac MRI
UKBB [35] - 31350 25080 6270
MESA [57] - 298 238 60
ACDC [4] - 300 240 60
MSCMR [59] - 300 240 60
Combined - 32248 25798 6450
Appendix B More Implementation Details
B.1 Data Preprocessing
Brain MRI data. Following [24, 29], we use 90 and 80 middle axial slices for BraTS and IXI datasets, respectively. These slices are further cropped to a size of $192\times 192$ from the central region. Ultimately, all volumes are resized to a fixed size of $192\times 192\times 64$ to serve as model input.
Cardiac MRI data. All slices within each cardiac MRI volume are used and cropped to $192\times 192$ from the central region. Each volume is then resized to a fixed size of $192\times 192\times 32$ for training and inference.
For all datasets, we apply intensity normalization by linearly scaling voxel intensities between the 0.5th and 99.5th percentiles to the range $[0,1]$. The data augmentations we employ include random spatial cropping, rotation, flipping, scaling, and shifting.
Table 9: Detailed architecture of each prompt encoder and projection head in the pretext task.
Prompt Encoder $\mathcal{F}_{1}$ $\mathcal{F}_{2}$ $\mathcal{F}_{3}$
Architecture
3D Conv
(in 8, out 256)
3D BatchNorm
ReLU
3D Conv
(in 256, out 512)
3D BatchNorm
ReLU
3D Adaptive Avg Pool
Linear (512, 1024)
ReLU
Linear (1024, 512)
3D Conv
(in 8, out 256)
3D BatchNorm
ReLU
3D Conv
(in 256, out 512)
3D BatchNorm
ReLU
3D Adaptive Avg Pool
Linear (512, 1024)
ReLU
Linear (1024, 512)
3D Conv
(in 8, out 256)
3D BatchNorm
ReLU
3D Conv
(in 256, out 512)
3D BatchNorm
ReLU
3D Adaptive Avg Pool
Linear (512, 1024)
ReLU
Linear (1024, 512)
Projection Head $\mathcal{H}_{1}$ $\mathcal{H}_{2}$ $\mathcal{H}_{3}$
Architecture
SiLU
Linear (512, m1m-1)
SiLU
Linear (512, $m$)
SiLU
Linear (512, 128)
B.2 Architecture of Prompt Encoders and Projection Heads
We devise three light weight prompt encoders to generate completeness-aware prompt tokens in our CoPeVAE, followed by three projection heads for each pretext task. The detailed architecture of each prompt encoder and projection head is illustrated in Table 9.
Notably, in our framework, binary mask codes are only used in the pretraining of CoPeVAE, where we synthetically remove modalities/slices and supervise the pretext tasks with known missing patterns. Once CoPeVAE is trained, we freeze its parameters and use it as a completeness-aware tokenizer: given any incomplete MRI with an arbitrary missing pattern, CoPeVAE directly infers the corresponding completeness prompts $\mathbf{p}=\mathbf{p}^{d}\,\|\,\mathbf{p}^{p}\,\|\,\mathbf{p}^{s}$ from the observed data, without requiring explicit mask codes as input. During both diffusion model training and inference, the diffusion backbone receives only the latent representations and these learned prompts. The original binary masks that were used to generate synthetic missingness are not provided to the diffusion model. In this sense, our synthesis process is mask-free: CoPeDiT no longer depends on hand-crafted or externally supplied mask codes at the generation stage, but instead relies entirely on the learned completeness prompts produced by the frozen CoPeVAE.
B.3 Hyperparameter Setups
CoPeVAE. The detailed hyperparameter setup of CoPeVAE is provided in Table 10. Built upon VQVAE [47] and VQGAN [10], our model employs a codebook containing 8192 codes with the latent dimensionality of 8. The values of $\tau$ and $\lambda$ are empirically set, as they have only a slight impact on model performance. The Adam optimizer is applied with a warmup cosine learning rate schedule. The training steps of CoPeVAE-B and CoPeVAE-C are 400k and 100k, respectively.
Figure 9: Qualitative results of comparison on IXI dataset. The results are depicted in the axial (top), sagittal (middle) and coronal (bottom) views of the 3D MRI volume.
Table 10: Hyperparameter setup of CoPeVAE.
CoPeVAE-B CoPeVAE-C
Architecture
Input dim. $m\times 192\times 192\times 64$ $192\times 192\times 32$
Num. codebook 8192 8192
Latent dim. 8 8
Channels (256, 384, 512) (32, 64, 128)
Compression ratio (8, 8, 8) (8, 8, 8)
Prompt dim. 512 512
$\tau$ 0.2 0.2
$\lambda$ 0.01 0.01
Optimization
Batch size 8 64
Learning rate 1e-4 1e-4
Optimizer Adam Adam
$(\beta_{1},\beta_{2})$ (0.9, 0.999) (0.9, 0.999)
LR schedule Warmup cosine Warmup cosine
Training steps 400k 200k
Table 11: Hyperparameter setup of MDiT3D.
MDiT3D-B MDiT3D-C
Architecture
Input dim. $m\times 8\times 24\times 24\times 8$ $4\times 8\times 24\times 24$
Hidden dim. 768 576
Num. blocks 16 12
Num. heads 12 12
Patch size 2 1
Params (M) 173.3 33.0
Flops (G) 555.1 (BraTS) 424.5 (IXI) 104.0
Optimization
Batch size 32 64
Learning rate 5e-5 5e-5
Optimizer AdamW AdamW
$(\beta_{1},\beta_{2})$ (0.9, 0.999) (0.9, 0.999)
LR schedule Warmup cosine Warmup cosine
Training steps 100k 100k
EMA decay 0.9999 0.9999
Interpolants
Training objective x-prediction x-prediction
Noise schedule scaled-linear scaled-linear
Timesteps 500 500
Sampler DDIM DDIM
Sampling steps 200 250
MDiT3D. For MDiT3D, we carefully design the hyperparameters to balance dataset size and model capacity, as summarized in Table 11. Following DiT [34], we report the experimental results using the exponential moving average (EMA) with a decay rate of 0.9999. During training, we set the timestep to 500 with linearly scaled noise levels ranging from 0.0015 to 0.0195. MDiT3D is trained for 100k iterations using the AdamW optimizer and a warmup cosine learning rate schedule. During inference, the DDIM sampler is applied with sample steps of 200 and 250 for MDiT3D-B and MDiT3D-C, respectively.
In addition, we use mixed-precision training with a gradient clipping to accelerate training and save computational resources throughout all two-stage experiments.
Appendix C Additional Experimental Results
C.1 Qualitative Results
Fig. 9 provides the qualitative evaluation results of brain MRI synthesis on the IXI dataset. As shown, we can also observe that our CoPeDiT outperforms other baselines in preserving the intricate structures and texture information in the synthesized MRIs.
Figure 10: Sensitivity of CoPeDiT to prompt corruption on BraTS dataset.
Table 12: Experimental results of tumor segmentation experiments on BraTS dataset.
[…]
WT TC ET AVG
Missing 86.08 84.67 81.59 84.11
GAN-based Methods
MMGAN [43] 89.35 88.14 87.73 88.41
MMT [24] 90.43 88.37 86.92 88.57
Hyper-GAE [53] 88.72 86.54 85.37 86.88
Diffusion Model-based Methods
LDM [39] 87.86 85.91 84.19 85.99
ControlNet [58] 88.27 87.05 85.23 86.85
M2DN [29] 91.28 90.09 88.20 89.86
CoPeDiT 91.35 90.41 88.94 90.23
C.2 Tumor Segmentation
In this section, we evaluate the clinical utility of our synthesized MRIs through downstream tumor segmentation experiments on the BraTS dataset. Specifically, following [24, 29], we train a multi-modal U-Net [40] on the four MRI modalities, where one modality is withheld and replaced by the synthesized image generated by our CoPeDiT or competing baselines. Additionally, we implement a U-Net that takes only the three available modalities as input for comparison. We report the average Dice scores for whole tumor (WT), tumor core (TC), and enhancing tumor (ET), as summarized in Table 12. As shown, all methods improve the segmentation accuracy over the ’Missing’ baseline that only uses the three available modalities, confirming the benefit of recovering the missing contrast. Among them, CoPeDiT achieves the best performance on all three tumor subregions, leading to the highest average Dice of 90.23%. These results demonstrate that our synthesized MRIs are not only visually plausible but also provide more informative inputs for downstream clinical tasks such as tumor segmentation.
Table 13: Ablation study on the contribution of completeness perception prompt tokens on the IXI dataset.
IXI
1 2
[…]
w/o $\mathbf{p}^{d}$ 23.56 0.720 35.93 23.04 0.708 48.71
w/o $\mathbf{p}^{p}$ 22.81 0.697 50.70 21.99 0.692 63.16
w/o $\mathbf{p}^{s}$ 24.17 0.724 31.89 23.56 0.710 39.87
w/o Prompts 22.48 0.696 57.64 21.67 0.683 81.49
w/ Mask Codes 23.35 0.711 44.03 22.84 0.702 56.72
CoPeDiT 24.34 0.732 25.84 23.92 0.721 32.53
Table 14: Quantitative results on the IXI dataset by incorporating our completeness perception prompt tokens into baseline methods instead of mask codes.
IXI
[…]
MMT [24] 22.64 0.698 53.60
+ Prompts (ours) 23.19 (+0.55) 0.707 (+0.009) 46.13 (–7.47)
Hyper-GAE [53] 22.12 0.682 72.62
+ Prompts (ours) 22.46 (+0.34) 0.694 (+0.012) 59.43 (–13.19)
M2DN [29] 23.47 0.715 42.52
+ Prompts (ours) 23.84 (+0.37) 0.726 (+0.011) 33.79 (–8.73)
C.3 Prompt Accuracy Sensitivity
To assess how sensitive CoPeDiT is to the correctness of completeness prompts, we conduct a controlled prompt corruption study on the BraTS dataset under the number of missing modalities of 1. Starting from a trained CoPeDiT and CoPeVAE, we consider the degree prompt $\mathbf{p}^{d}$ (number/length of missing content), the position prompt $\mathbf{p}^{p}$ (location of missing regions), and the semantic prompt $\mathbf{p}^{s}$. For each prompt type, we define a corruption rate $r\in\{0,0.25,0.5,0.75,1.0\}$. Given a rate $r$, we randomly select an $r$-fraction of validation samples and replace the corresponding prompt with an incorrect one (sampled from other cases), while keeping all network weights and the other prompts unchanged. We also include an ’All’ setting where all three prompts are corrupted simultaneously. As illustrated in Fig. 10, increasing the corruption rate consistently degrades synthesis quality in terms of all three metrics. This demonstrates that CoPeDiT relies on the accuracy of completeness prompts. Among the three types, corrupting the position prompt $\mathbf{p}^{p}$ causes the largest synthesis performance drop, emphasizing that precise localization of missing regions is particularly critical. Corrupting the degree prompt $\mathbf{p}^{d}$ also noticeably harms performance, whereas the semantic prompt $\mathbf{p}^{s}$ is somewhat more robust but still shows a clear degradation trend. When all prompts are corrupted, performance deteriorates the most across all metrics, confirming that the three prompts provide complementary and influential guidance for high-fidelity synthesis.
C.4 Ablation Study
We further evaluate the effectiveness of our prompt token design on the IXI dataset. As shown in Table 13, the complete set of prompt tokens yields the best performance, outperforming both conventional mask codes and all individual prompt tokens. Furthermore, we apply our learned prompt tokens to other baseline models originally using mask codes. As illustrated in Table 14, our prompts also lead to consistent performance gains across all baselines. In summary, the effectiveness of our prompt learning scheme is validated on the IXI dataset through the additional evaluations presented above.

### v2

@@
 Towards this goal, we propose CoPeDiT, a general-purpose latent diffusion model equipped with completeness perception for unified synthesis of 3D MRIs.
-Specifically, we incorporate dedicated pretext tasks into our tokenizer, CoPeVAE, empowering it to learn completeness-aware discriminative prompts, and design MDiT3D, a specialized diffusion transformer architecture for 3D MRI synthesis, that effectively uses the learned prompts as guidance to enhance semantic consistency in 3D space.
-Comprehensive evaluations on three large-scale MRI datasets demonstrate that CoPeDiT significantly outperforms state-of-the-art methods, achieving superior robustness, generalizability, and flexibility.
-The code is available at https://github.com/JK-Liu7/CoPeDiT.
+Specifically, we incorporate dedicated pretext tasks into our tokenizer, CoPeVAE, empowering it to learn completeness-aware discriminative prompts, and design MDiT3D, a specialized diffusion transformer architecture for 3D MRI synthesis that effectively uses the learned prompts as guidance to enhance semantic consistency in 3D space.
+Comprehensive evaluations on three large-scale MRI datasets demonstrate that CoPeDiT significantly outperforms state-of-the-art methods, achieving superior robustness and yielding high-fidelity, structurally consistent synthesis across diverse missing patterns.
 1 Introduction
-Magnetic resonance imaging (MRI) is widely used in clinical medicine due to its non-invasive nature and ability to visualize detailed tissue properties [8].
-For instance, multi-modal brain MRIs provide complementary information regarding brain anatomy and pathology [27], while volumetric cardiac MRIs reveal detailed cardiac structure and tissue characteristics [7].
-However, brain and cardiac MRIs often suffer from missing data in real-world clinical settings [50].
-For brain MRIs, certain modalities may be unavailable, while for cardiac MRIs, some slices may be missing, which are commonly caused by limited scanning time, image corruption, and variations in acquisition protocols [33].
-Figure 1: The motivation of CoPeDiT.
-(a) Comparison of completeness perception between CoPeDiT and prior methods.
-(b) Prompt tokens offer more effective guidance than binary mask codes.
-(c) CoPeVAE enables more discriminative latent representations.
-(d) Prompt tokens yield more semantically consistent attention maps in MDiT3D blocks, highlighting stronger interactions between similar modalities (e.g., T1-T1ce, T2-FLAIR).
-To address this, generative models have been developed to infer missing data from observed inputs [19, 11].
-Existing paradigms rely on auxiliary embeddings, e.g., binary mask codes, as prior knowledge to encode missing patterns (e.g., severity, type, and position) [24, 14, 21, 6, 29].
-Nonetheless, this approach has several inherent limitations that hinder its practical applicability.
-First, in real-world deployments, input images may originate from different hospitals, scanners, or time points.
-As missing patterns are intrinsically unpredictable, providing manually defined masks in advance is impractical [49, 41].
-Second, this over-reliance on predefined masks causes generative models to be unaware of the missing state and unable to distinguish between different modalities or spatial characteristics, thereby reducing robustness to variable and unseen incomplete patterns [22, 20, 51].
-Models conditioned on fixed mask representations often suffer from limited generalizability, exhibiting noticeable drop in synthesis quality when encountering unfamiliar missing sites [13, 2, 32].
-Finally, binary mask codes lack semantic richness and adaptability, leading to a rigid guidance that weakens spatial alignment and semantic coherence in the synthesized output [37, 17, 44].
-Intuitively, generative models should be capable of inferring and detecting the incomplete state spontaneously, rather than relying on externally manual guidance [17, 12].
-Motivated by this, we pose a central question: ‘Can we empower the model with the ability to perceive missing states on its own?’ In light of this, we exploit a long neglected and underexplored property of generative models in medical imaging, i.e., completeness perception, to enhance flexibility and generalizability under arbitrary missing MRI conditions, as illustrated in Fig.
-1 (a).
+Magnetic resonance imaging (MRI) provides crucial anatomical and pathological insights, particularly through multi-modal brain and volumetric cardiac scans Dickinson et al.
+(2013); Lustig et al.
+(2007); Dayarathna et al.
+(2024); Manna et al.
+(2024).
+However, real-world clinical MRIs frequently suffer from missing data, including absent brain modalities and missing cardiac slices, due to limited scan times, image corruption, or protocol variations Wang et al.
+(2025b); Paproki et al.
+(2024); Zhao & Shen (2025).
+Figure 1: Overview and advantages of CoPeDiT.
+(a) We shift the paradigm from explicit mask-dependent guidance to autonomous completeness perception via prompt generation.
+This self-perceptive mechanism offers significant advantages: (b) quantitative plug-and-play performance gains across existing baselines; (c) substantially more discriminative latent representations compared to training without pretext tasks; and (d) enhanced semantic attention alignment between correlated modalities (e.g., T1-T1ce, T2-FLAIR) compared to traditional binary mask codes.
+To address this, generative models have been developed to infer missing data from observed inputs Ibrahim et al.
+(2025); Ferreira et al.
+(2024).
+Existing paradigms rely on auxiliary embeddings, e.g., binary mask codes, as prior knowledge to encode missing patterns (e.g., severity, type, and position) Liu et al.
+(2023); Hao et al.
+(2024); Kim & Park (2024); Cho et al.
+(2024); Meng et al.
+(2024).
+Nonetheless, these hand-crafted masks only indicate missing locations, without adequately characterizing the actual incomplete state of the input.
+This causes three practical limitations.
+First, because missing patterns vary across hospitals, scanners, and acquisition settings, enumerating them with predefined masks is unrealistic in real deployments Wang et al.
+(2023); Rui et al.
+(2025).
+Second, the resulting condition is insensitive to modality-specific and spatially varying context, making models less robust to unseen incomplete patterns and prone to degraded generalization Lee et al.
+(2023); Ke et al.
+(2025); Wenderoth et al.
+(2025); Hamamci et al.
+(2024); Azad et al.
+(2025); Pan et al.
+(2025).
+Third, because binary masks carry limited semantics, they provide rigid and insufficiently informative guidance, which can weaken spatial alignment and semantic consistency during synthesis Qiu et al.
+(2023); Hu et al.
+(2023); Shin et al.
+(2025); Wu et al.
+(2025).
+Intuitively, generative models should be capable of inferring and detecting the incomplete state spontaneously, rather than relying on externally provided manual guidance Hu et al.
+(2023); Graikos et al.
+(2024).
+Motivated by this, we pose a central question: Can we empower the model with the ability to perceive missing states on its own?
+In light of this, we exploit an underexplored property of generative models in medical imaging, i.e., completeness perception, to enhance flexibility and generalizability under arbitrary missing MRI conditions, as illustrated in Fig.
+1.
 Our fundamental insight is to enable the generative model to recognize the fine-grained incomplete state information in a self-perceptive manner, and to leverage this understanding as internal prompts to guide the generation process.
 We hypothesize that, for diffusion models, such self-guided prompts may serve as an effective alternative to manually defined masks, and potentially offer even stronger guidance signals (Fig.
-1 (b)).
-The main reason lies in the fact that this self-perceptive strategy encourages the model to learn both global and local anatomical structures and lesion patterns at coarse and fine levels, thus enabling more semantically coherent generation of the missing MRI regions during synthesis (Fig.
-1 (c)(d)) [23, 18].
+1).
+The main reason lies in the fact that this self-perceptive strategy encourages the model to learn both global and local anatomical structures and lesion patterns at coarse and fine levels, thus enabling more semantically coherent generation of the missing MRI regions during synthesis (Figs.
+11Liang et al.
+(2022).
 Driven by our motivation, we propose CoPeDiT, a 3D latent diffusion model (LDM) framework for unified 3D MRI synthesis.
+Here, our “unified” denotes a common formulation, training paradigm, and conditioning strategy, with task-specific adaptations for different synthesis settings.
 Technically, our framework builds on two core components: (i) Unlike prior approaches that require explicit missing indicators, a novel tokenizer with a completeness perception function, CoPeVAE, is proposed to autonomously assess the integrity of modalities or volumes through tailored self-supervised pretext tasks.
 By detecting anatomical structures and variations in lesion patterns, CoPeVAE develops a comprehensive understanding of 3D MRIs.
-This enables CoPeDiT to eliminate the need for manual intervention with flexible adaptability, enhancing the method’s autonomy and improves its feasibility for real-world clinical deployment with diverse missing patterns.
-(ii) A novel diffusion transformer architecture, MDiT3D, is developed to unlock the potential of Diffusion Transformers (DiT) [34] for medical image synthesis, offering a new paradigm for generating high-quality volumetric data.
-MDiT3D leverages tokenization and attention mechanisms to model long-range and irregular contextual dependencies in high-dimensional, structurally complex 3D MRIs.
-It supports conditional generation guided by completeness-aware prompts, which provide semantic-meaningful guidance on the synthesis process.
-The learned prompt tokens comprehensively absorb information from available modalities or slices during the diffusion process and guide the generation of the missing ones.
-Incorporating the above two innovations, our architecture not only enables adaptive self-guidance synthesis in a mask-free manner which exhibits strong robustness and generalizability, but also demonstrates improved structural coherence and enhanced preservation of fine-grained anatomical details.
+This enables CoPeDiT to eliminate the need for manual intervention with flexible adaptability, enhancing the method’s autonomy and improving its feasibility for real-world clinical deployment with diverse missing patterns.
+(ii) A task-specific diffusion transformer instantiation, MDiT3D, is developed as a dependency-aligned conditioning interface for completeness-aware prompts in 3D MRI synthesis.
+Rather than introducing a fundamentally new DiT paradigm, MDiT3D adapts tokenization, attention, and prompt injection to the long-range, anisotropic, and irregular structural dependencies of volumetric MRIs.
+This design provides the proper pathway for our learned prompts to influence the generative process, allowing the prompt tokens to propagate observed modality- or slice-specific cues along anatomically meaningful dependencies during diffusion.
+As a result, MDiT3D enables more reliable synthesis of missing modalities or slices while better preserving structural consistency in high-dimensional 3D data.
+Incorporating the above two innovations, our architecture not only enables adaptive self-guidance synthesis, but also demonstrates improved structural coherence and enhanced preservation of fine-grained anatomical details.
 Our main contributions are summarized as follows:
-  • We propose a unified framework, dubbed CoPeDiT, for both 3D brain and cardiac missing MRI synthesis under arbitrary incomplete scenarios, without the need for explicit external indicators as guidance.
+  • We propose a unified formulation with task-specific instantiations, dubbed CoPeDiT, for both 3D brain and cardiac missing MRI synthesis under arbitrary incomplete scenarios, without the need for explicit external indicators as guidance.
   • We empower our tokenizer, CoPeVAE, with a strong capacity to perceive completeness by seamlessly integrating carefully designed pretext tasks, enabling the model to recognize missing states and learn informative, self-guided prompts.
-  • We present MDiT3D, a novel latent diffusion transformer for 3D MRI synthesis, featuring a dedicated architecture for capturing modality/spatial relationships and leveraging generated prompts as conditional guidance.
-  • Extensive experiments on three datasets demonstrate that our model surpasses state-of-the-art (SOTA) methods, achieving generalizability, robustness, and real-world clinical applicability.
+  • We present MDiT3D, a tailored diffusion transformer for 3D MRI synthesis, designed as a dependency-aligned conditioning interface for our completeness-aware prompts to enable structurally meaningful prompt-guided generation.
+  • Extensive experiments on three datasets demonstrate that our model surpasses state-of-the-art (SOTA) methods, achieving robustness and real-world clinical applicability.
 2 Related Work
+Medical Image Generation.
+Medical image generation has been widely studied for data augmentation Hamamci et al.
+(2024); Zhao et al.
+(2025), reconstruction Yu et al.
+(2025); Xiao et al.
+(2025); Liu et al.
+(2024), and image completion Rassmann et al.
+(2025); Liu et al.
+(2023); Song et al.
+(2026) in clinical imaging workflows.
+Earlier methods based on Generative Adversarial Networks (GANs) improved realism but frequently faced limitations in training stability and mode collapse, restricting their fidelity, especially when scaling to complex, high-dimensional 3D volumetric data Cao et al.
+(2023); Shao et al.
+(2025); Weng et al.
+(2024).
+To overcome these bottlenecks, diffusion-based models have emerged as a robust alternative.
+By offering stable training dynamics and superior mode coverage, they have achieved competitive performance in medical image synthesis and are increasingly favored in conditional settings guided by partial observations, anatomical priors, or auxiliary modalities Wang et al.
+(2025a); Zhao et al.
+(2025); Zhang et al.
+(2026); Guo et al.
+(2025); Yeganeh et al.
+(2025).
+MRI Synthesis.
+Recent unified MRI synthesis methods predominantly utilize generative models, ranging from GANs Xia et al.
+(2021); Zhang et al.
+(2024; 2019a; 2019b); Yang et al.
+(2023); Sharma & Hamarneh (2020) and Transformers Liu et al.
+(2023; 2021; 2025) to advanced diffusion models Qiu et al.
+(2025); Yeganeh et al.
+(2025); Meng et al.
+(2024); Song et al.
+(2026).
+While these approaches successfully capture complex inter-modal dependencies to impute missing data, they inherently rely on externally provided masks to explicitly encode missing patterns in randomly incomplete scenarios Liu et al.
+(2023); Meng et al.
+(2024); Wang et al.
+(2023); Azad et al.
+(2025).
+This manual guidance is often rigid and lacks informative semantic details about the actual incomplete state of the input.
+Furthermore, because missing patterns vary widely across different clinical environments, requiring predefined masks limits practical deployment.
+In contrast to these mask-dependent paradigms, CoPeDiT explores the self-perceptive capability of generative models to autonomously recognize data completeness.
+By learning to infer missing states internally, our approach eliminates the need for manual intervention, enabling more flexible and high-fidelity MRI synthesis.
+Diffusion Models.
+Diffusion models, particularly LDMs Rombach et al.
+(2022), have demonstrated remarkable capabilities across various vision tasks Ho et al.
+(2020); Lu et al.
+(2024); Ma et al.
+(2025); Yao et al.
+(2025).
+Recently, DiTs Peebles & Xie (2023) have emerged as powerful alternatives to traditional U-Net Ronneberger et al.
+(2015) backbones, achieving competitive performance in natural image synthesis.
+However, most existing medical image generation approaches still predominantly rely on U-Net architectures, leaving the potential of transformer-based LDMs largely underexplored in this domain Wang et al.
+(2025a); Nazir et al.
+(2025).
+To bridge this gap, we introduce MDiT3D, which replaces the U-Net backbone with a diffusion transformer.
+By incorporating task-specific architectural modifications, MDiT3D effectively captures the complex, long-range dependencies inherent in high-dimensional 3D MRIs.
+3 Methodology
 Figure 2: The overview framework of CoPeVAE. We implement two variants, CoPeVAE-B and CoPeVAE-C, with slight architectural modifications for brain and cardiac MRI synthesis tasks, respectively.
-MRI Synthesis.
-Recent efforts focus on unified MRI synthesis using generative models such as Generative Adversarial Networks (GAN) and diffusion models [52, 56, 1, 36].
-For example, MMGAN [43] employs a single network to handle diverse missing configurations, while Hyper-GAE [53] introduces a shared hyper-encoder and decoder architecture with a graph-based fusion module for reconstructing missing brain MRIs.
-MMT [24] proposes to leverage Swin Transformers [25] to capture inter-modal dependencies.
-M2DN [29] presents a diffusion-based method that jointly models all modalities using a multi-input multi-output framework.
-Despite these advances, most existing methods still rely on externally provided masks to encode missing patterns in randomly incomplete cases.
-In contrast, CoPeDiT explores self-perceptive capability of generative models to autonomously recognize data completeness, enabling more flexible and high-fidelity MRI synthesis.
-Diffusion Models.
-Diffusion models have shown strong generative capacities across various computer vision tasks [26, 28, 55].
-Denoising Diffusion Probabilistic Models [16] generate images by gradually adding Gaussian noise and learning to reverse the process.
-LDM [39] improve efficiency by operating in a compressed latent space using autoencoders.
-Recently, Diffusion transformers [34] have been effective alternatives to U-Net [40] in LDMs, achieving competitive performance in natural image generation.
-However, most prior medical image generation approaches still rely on U-Net-based architectures as the diffusion backbone, leaving the potential of transformer-based diffusion models underexplored in this domain [48, 31].
-In this paper, we devise MDiT3D, by replacing the U-Net backbone with a diffusion transformer and incorporating task-specific architectural modifications, MDiT3D is capable of generalizing to 3D MRI data and capturing long-range dependencies of high-dimensional 3D MRIs.
-Figure 3: The overview framework of MDiT3D. PE and RoPE denote positional embeddings and rotary position embeddings [45, 54], respectively.
-3 Methodology
 3.1 Notations
-We consider two 3D MRI synthesis tasks: brain missing modality synthesis and cardiac missing slice synthesis.
-To unify the formulation, let $\mathcal{M}=\{\mathbf{x}^{i}\}_{i=1}^{m}$ denote a complete MRI sample, where $m$ is the total number of modalities or slices.
-The available and missing subsets are denoted as $\mathcal{M}^{S}=\{\mathbf{x}^{s_{i}}\}_{i=1}^{s}$ and $\mathcal{M}^{C}=\{\mathbf{x}^{c_{i}}\}_{i=1}^{c}$, respectively, with $m=s+c$.
-(1) Brain missing modality synthesis: $\mathcal{M}$ represents all brain MRIs with $m$ modalities, and the task is to infer $\mathcal{M}^{C}$ from $\mathcal{M}^{S}$.
-(2) Cardiac missing slice synthesis: $\mathcal{M}$ corresponds to a volume with $m$ slices, where missing slices $\mathcal{M}^{C}$ are consecutive, non-overlapping with $\mathcal{M}^{S}$, and are to be reconstructed from the available ones.
-Notably, our experimental setup considers randomly generated incomplete cases from all the possible combinations as input, with a varying number and length of missing modalities or slices, aiming to mimic real-world clinical environments.
+To unify the formulation of brain missing-modality synthesis and cardiac missing-slice synthesis, let $\mathcal{M}=\{\mathbf{x}^{i}\}_{i=1}^{m}$ denote a complete 3D MRI sample comprising $m$ elements (either modalities or slices).
+We partition $\mathcal{M}$ into an available subset $\mathcal{M}^{S}=\{\mathbf{x}^{s_{i}}\}_{i=1}^{s}$ and a missing subset $\mathcal{M}^{C}=\{\mathbf{x}^{c_{i}}\}_{i=1}^{c}$, where $m=s+c$.
+The objective is to synthesize $\mathcal{M}^{C}$ from $\mathcal{M}^{S}$.
+Specifically, for brain MRIs, $\mathcal{M}^{C}$ contains randomly missing modalities; for cardiac MRIs, $\mathcal{M}^{C}$ consists of consecutive missing slices.
+To mimic real-world clinical environments, we evaluate randomly generated incomplete cases with varying missing counts and lengths.
 3.2 Stage I: Completeness Perception Tokenizer
@@
 2) is that detecting the completeness of high-resolution MRI data enforces the model to perceive both global anatomy and local lesion patterns, thereby producing high-quality prompts as diffusion guidance.
-Building upon VQGAN [47, 10], we deploy a 3D autoencoder jointly trained with self-supervised pretext tasks.
+Building upon VQGAN van den Oord et al.
+(2017); Esser et al.
+(2021), we deploy a 3D autoencoder jointly trained with self-supervised pretext tasks.
 Each task employs a prompt encoder, denoted as $\mathcal{F}_{1},\mathcal{F}_{2}$ and $\mathcal{F}_{3}$, to transform latent tokens (learned by the encoder $\mathcal{E}$) into low-dimensional prompt tokens.
@@
 Missing Number/Length Detection.
-Equipped with global contextual awareness, the tokenizer is capable of answering the question that how many modalities/slices are missing in the incomplete input.
+Equipped with global contextual awareness, the tokenizer is capable of determining how many modalities/slices are missing in the incomplete input.
 This task aims to enable the tokenizer to identify the severity of incompleteness by perceiving the global context of MRIs, thereby learning modality/spatial attributes in a coarse-grained manner.
-We formulate this task as an (m1)(m-1)-class classification task and define the loss using the cross-entropy as
-$\small\mathcal{L}_{d}=\mathcal{L}_{\text{cls}}(\mathcal{H}_{1}(\mathcal{F}_{1}(\mathbf{z}^{s}))_{c}).$ (1)
+We formulate this task as an (m1)(m-1)-class classification task and define the loss using cross-entropy loss as follows:
+$\mathcal{L}_{d}=\mathcal{L}_{\text{cls}}(\mathcal{H}_{1}(\mathcal{F}_{1}(\mathbf{z}^{s}))_{c}).$ (1)
 where $\mathcal{F}_{1}$ and $\mathcal{H}_{1}$ represent the prompt encoder and the projection head for missing detection, respectively.
-The learned prompt tokens $\mathbf{p}^{d}=\mathcal{F}_{1}(\mathbf{z}^{s})$ are rich in information about the severity of missing state.
+The learned prompt tokens $\mathbf{p}^{d}=\mathcal{F}_{1}(\mathbf{z}^{s})$ are rich in information about the severity of the missing state.
+Figure 3: Architecture of the MDiT3D framework.
+We implement two variants with task-specific alternating blocks (Spatial/Modal for MDiT3D-B and Planar/Spatial for MDiT3D-C), using dynamic feature reshaping to accurately model respective anatomical dependencies.
+Learned prompts are injected via adaLN as conditional guidance.
+PE: Positional Embeddings; RoPE: Rotary Position Embeddings Su et al.
+(2024); Yang et al.
+(2025).
 Task 2.
@@
 This task is formulated as an $m$-class classification problem and is also optimized using the cross-entropy loss, defined as follows:
-$\small\mathcal{L}_{p}=\mathcal{L}_{\text{cls}}(\mathcal{H}_{2}(\mathcal{F}_{2}(\mathbf{z}^{s}))_{I}).$ (2)
+$\mathcal{L}_{p}=\mathcal{L}_{\text{cls}}(\mathcal{H}_{2}(\mathcal{F}_{2}(\mathbf{z}^{s}))_{I}).$ (2)
 where $I$ denotes the index of missing type/position.
@@
 Missing Modality/Slice Assessment.
-Motivated by the observation that modalities or slices from the same scan share more similar anatomical and textural context than those from different scans, we adopt an inter-modal/inter-slice contrastive learning scheme [38] to serve as a missing data estimator.
-Specifically, we take the incomplete latent tokens $\mathbf{z}^{s}$ as anchor, the corresponding missing latent $\mathbf{z}^{c}$ from the same subject as positives, and latent tokens $\mathbf{z}^{c}_{-}$ from different subjects as negatives.
+Motivated by the observation that modalities or slices from the same scan share more similar anatomical and textural context than those from different scans, we adopt an inter-modal/inter-slice contrastive learning scheme Radford et al.
+(2021) to serve as a missing data estimator.
+Specifically, we take the incomplete latent tokens $\mathbf{z}^{s}$ as the anchor, the corresponding missing latent $\mathbf{z}^{c}$ from the same subject as positives, and latent tokens $\mathbf{z}^{c}_{-}$ from different subjects as negatives.
 The contrastive loss is defined as follows:
-$\mathcal{L}_{c}=-\log\frac{\varphi\left(\mathcal{H}_{3}(\mathbf{p}^{s}),\mathcal{H}_{3}(\mathbf{p}^{c})\right)}{\varphi\left(\mathcal{H}_{3}(\mathbf{p}^{s}),\mathcal{H}_{3}(\mathbf{p}^{c})\right)+\sum\varphi\left(\mathcal{H}_{3}(\mathbf{p}^{s}),\mathcal{H}_{3}(\mathbf{p}^{c}_{-})\right)}.$
+$\mathcal{L}_{c}=-\log\frac{\varphi\left(\mathcal{H}_{3}(\mathbf{p}^{s}),\mathcal{H}_{3}(\mathbf{p}^{c})\right)}{\varphi\left(\mathcal{H}_{3}(\mathbf{p}^{s}),\mathcal{H}_{3}(\mathbf{p}^{c})\right)+\sum\varphi\left(\mathcal{H}_{3}(\mathbf{p}^{s}),\mathcal{H}_{3}(\mathbf{p}^{c}_{-})\right)}.$ (3)
 where $\mathbf{p}^{s}=\mathcal{F}_{3}(\mathbf{z}^{s})$, $\quad\mathbf{p}^{c}=\mathcal{F}_{3}(\mathbf{z}^{c})$, $\varphi(a,b)=\exp(a\cdot b/\tau)$, and $\tau$ is the temperature parameter.
-This contrastive learning scheme encourages the model to focus on inter-modal/slice contextual differences, thus improving anatomical coherence and fine-grained detail preservation.
+This contrastive learning scheme encourages the model to focus on inter-modal/slice contextual differences, improving anatomical coherence and fine-grained detail preservation.
 The overall loss of CoPeVAE is formulated as
-$\small\mathcal{L}_{\text{tok}}=\mathcal{L}_{\text{rec}}+\lambda(\mathcal{L}_{d}+\mathcal{L}_{p}+\mathcal{L}_{c}).$ (4)
-where $\lambda$ is the weighting coefficient, and $\mathcal{L}_{\text{rec}}$ denotes the reconstruction loss that consists of a pixel-wise L1 loss, vector quantization loss, adversarial loss, and perceptual loss.
-Table 1: Quantitative results for multi-modal brain MRI synthesis on BraTS and IXI datasets. The numbers of the second row denote the number of missing modalities. The best performances are bolded.
-BraTS IXI
-1 2 3 1 2
+$\mathcal{L}_{\text{tok}}=\mathcal{L}_{\text{rec}}+\lambda(\mathcal{L}_{d}+\mathcal{L}_{p}+\mathcal{L}_{c}).$ (4)
+where $\lambda$ is the weighting coefficient, and $\mathcal{L}_{\text{rec}}$ is the reconstruction loss containing a L1 loss, vector quantization loss, adversarial loss, and perceptual loss.
+Table 1: Quantitative results for multi-modal brain MRI synthesis on the BraTS dataset. The numbers in the first row denote the number of missing modalities.
+1 2 3
 […]
 GAN-based Methods
-MMGAN [43] 24.71 $\pm$1.57 0.817 $\pm$0.027 27.94 $\pm$1.05 24.38 $\pm$1.74 0.806 $\pm$0.023 32.48 $\pm$1.66 24.06 $\pm$1.92 0.794 $\pm$0.031 39.37 $\pm$2.10 22.29 $\pm$1.35 0.684 $\pm$0.015 70.91 $\pm$3.64 21.13 $\pm$1.49 0.668 $\pm$0.019 93.57 $\pm$3.42
-MMT [24] 25.19 $\pm$1.41 0.824 $\pm$0.017 24.53 $\pm$1.33 24.50 $\pm$1.55 0.811 $\pm$0.020 29.57 $\pm$1.82 23.92 $\pm$1.53 0.801 $\pm$0.021 39.66 $\pm$1.95 22.64 $\pm$1.49 0.698 $\pm$0.013 53.60 $\pm$3.14 21.82 $\pm$1.60 0.687 $\pm$0.019 72.24 $\pm$2.98
-Hyper-GAE [53] 24.65 $\pm$1.62 0.813 $\pm$0.022 28.97 $\pm$1.57 24.42 $\pm$1.74 0.808 $\pm$0.024 33.52 $\pm$1.78 23.86 $\pm$1.88 0.788 $\pm$0.029 41.79 $\pm$2.21 22.12 $\pm$1.33 0.682 $\pm$0.017 72.62 $\pm$3.78 20.91 $\pm$1.45 0.662 $\pm$0.021 98.79 $\pm$4.05
+MMGAN Sharma & Hamarneh (2020) 24.71 $\pm$1.57 0.817 $\pm$0.027 27.94 489.86 24.38 $\pm$1.74 0.806 $\pm$0.023 32.48 726.93 24.06 $\pm$1.92 0.794 $\pm$0.031 39.37 898.17
+MMT Liu et al. (2023) 25.19 $\pm$1.41 0.824 $\pm$0.017 24.53 527.06 24.50 $\pm$1.55 0.811 $\pm$0.020 29.57 686.52 23.92 $\pm$1.53 0.801 $\pm$0.021 39.66 841.46
+Hyper-GAE Yang et al. (2023) 24.65 $\pm$1.62 0.813 $\pm$0.022 28.97 609.65 24.42 $\pm$1.74 0.808 $\pm$0.024 33.52 815.92 23.86 $\pm$1.88 0.788 $\pm$0.029 41.79 948.27
 Diffusion Model-based Methods
-LDM [39] 23.84 $\pm$1.52 0.805 $\pm$0.019 36.47 $\pm$1.54 23.12 $\pm$1.64 0.798 $\pm$0.021 45.93 $\pm$2.41 22.65 $\pm$1.77 0.791 $\pm$0.025 52.50 $\pm$2.07 21.36 $\pm$1.28 0.679 $\pm$0.016 86.62 $\pm$3.86 20.94 $\pm$1.53 0.654 $\pm$0.020 112.57 $\pm$5.02
-ControlNet [58] 23.98 $\pm$1.49 0.808 $\pm$0.018 34.09 $\pm$1.93 23.34 $\pm$1.60 0.801 $\pm$0.020 41.28 $\pm$2.49 22.85 $\pm$1.68 0.795 $\pm$0.022 48.07 $\pm$1.99 21.83 $\pm$1.51 0.681 $\pm$0.015 81.26 $\pm$3.85 21.07 $\pm$1.50 0.661 $\pm$0.019 103.77 $\pm$5.17
-M2DN [29] 26.45 $\pm$1.36 0.830 $\pm$0.016 21.29 $\pm$1.18 25.87 $\pm$1.48 0.820 $\pm$0.017 27.36 $\pm$1.62 25.21 $\pm$1.59 0.809 $\pm$0.024 32.40 $\pm$1.93 23.47 $\pm$1.43 0.715 $\pm$0.014 42.52 $\pm$3.08 22.81 $\pm$1.56 0.702 $\pm$0.018 55.64 $\pm$3.62
-CoPeDiT 28.26 $\pm$1.24 0.842 $\pm$0.019 12.67 $\pm$0.98 28.13 $\pm$1.49 0.831 $\pm$0.021 13.25 $\pm$1.45 27.91 $\pm$1.41 0.822 $\pm$0.023 14.89 $\pm$1.61 24.34 $\pm$1.21 0.732 $\pm$0.016 25.84 $\pm$2.58 23.92 $\pm$1.53 0.721 $\pm$0.020 32.53 $\pm$3.16
+LDM Rombach et al. (2022) 23.84 $\pm$1.52 0.805 $\pm$0.019 36.47 740.26 23.12 $\pm$1.64 0.798 $\pm$0.021 45.93 897.83 22.65 $\pm$1.77 0.791 $\pm$0.025 52.50 1161.21
+ControlNet Zhang et al. (2023) 23.98 $\pm$1.49 0.808 $\pm$0.018 34.09 806.82 23.34 $\pm$1.60 0.801 $\pm$0.020 41.28 986.30 22.85 $\pm$1.68 0.795 $\pm$0.022 48.07 1074.23
+M2DN Meng et al. (2024) 26.45 $\pm$1.36 0.830 $\pm$0.016 21.29 376.53 25.87 $\pm$1.48 0.820 $\pm$0.017 27.36 467.66 25.21 $\pm$1.59 0.809 $\pm$0.024 32.40 553.16
+DiffM4RI Ye et al. (2026) 26.07 $\pm$1.57 0.824 $\pm$0.024 25.03 449.17 25.49 $\pm$1.41 0.813 $\pm$0.025 32.59 591.35 25.08 $\pm$1.75 0.806 $\pm$0.029 35.91 695.06
+CoPeDiT 28.26 $\pm$1.24 0.842 $\pm$0.019 12.67 254.71 28.13 $\pm$1.49 0.831 $\pm$0.021 13.25 287.58 27.91 $\pm$1.41 0.822 $\pm$0.023 14.89 323.19
+Table 2: Quantitative results for multi-modal brain MRI synthesis on the IXI dataset. The numbers in the first row denote the number of missing modalities.
+1 2
+[…]
+GAN-based Methods
+MMGAN Sharma & Hamarneh (2020) 22.29 $\pm$1.35 0.684 $\pm$0.015 70.91 1447.83 21.13 $\pm$1.49 0.668 $\pm$0.019 93.57 1787.39
+MMT Liu et al. (2023) 22.64 $\pm$1.49 0.698 $\pm$0.013 53.60 1329.25 21.82 $\pm$1.60 0.687 $\pm$0.019 72.24 1562.44
+Hyper-GAE Yang et al. (2023) 22.12 $\pm$1.33 0.682 $\pm$0.017 72.62 1520.49 20.91 $\pm$1.45 0.662 $\pm$0.021 98.79 1712.90
+Diffusion Model-based Methods
+LDM Rombach et al. (2022) 21.36 $\pm$1.28 0.679 $\pm$0.016 86.62 1884.64 20.94 $\pm$1.53 0.654 $\pm$0.020 112.57 2314.39
+ControlNet Zhang et al. (2023) 21.83 $\pm$1.51 0.681 $\pm$0.015 81.26 1971.40 21.07 $\pm$1.50 0.661 $\pm$0.019 103.77 2386.04
+M2DN Meng et al. (2024) 23.47 $\pm$1.43 0.715 $\pm$0.014 42.52 845.29 22.81 $\pm$1.56 0.702 $\pm$0.018 55.64 1078.36
+DiffM4RI Ye et al. (2026) 23.71 $\pm$1.26 0.720 $\pm$0.018 39.87 715.63 22.87 $\pm$1.68 0.707 $\pm$0.022 51.93 964.28
+CoPeDiT 24.34 $\pm$1.21 0.732 $\pm$0.016 25.84 569.22 23.92 $\pm$1.53 0.721 $\pm$0.020 32.53 718.54
 3.3 Stage II: 3D MRI Diffusion Transformer
 We propose MDiT3D (see Fig.
-3), extending the model configuration of DiT [34] with several key design innovations to enable MDiT3D to adapt to 3D modalities and volumetric MRI.
-It takes the high-dimensional latent tokens $\mathbf{z}^{s}$ as input, along with the learned prompt tokens $\mathbf{p}=\mathbf{p}^{d}\,\|\,\mathbf{p}^{p}\,\|\,\mathbf{p}^{s}$ from our CoPeVAE tokenizer, which encode comprehensive missing state information and serve as conditional guidance for synthesizing.
-Without loss of generality, we use MDiT3D-B as a representative configuration to illustrate the architectural design.
-3D Patchify & 3D Positional Embeddings.
-We apply a 3D patchify operator with patch size $p\times p\times p$ to obtain $\mathbf{h}\in\mathbb{R}^{m\times T\times d}$ [30], where $m$ denotes the number of modality, and $d$ is the embedding dimension.
-Unlike the standard 2D positional embeddings [9] used in DiT, we leverage 3D frequency-based sine-cosine positional embeddings (3D PE) to better encode spatial positional information within the 3D MRI data.
+3), a task-specific diffusion transformer extending DiT Peebles & Xie (2023) for volumetric MRI synthesis.
+MDiT3D is conditionally guided by the available latents $\mathbf{z}^{s}$ alongside the concatenated completeness-aware prompts $\mathbf{p}=\mathbf{p}^{d}\,\|\,\mathbf{p}^{p}\,\|\,\mathbf{p}^{s}$.
+During inference, these prompts are autonomously extracted by the frozen CoPeVAE to provide semantic guidance: $\mathbf{p}^{d}$ calibrates global severity (how many), $\mathbf{p}^{p}$ localizes missing regions (where), and $\mathbf{p}^{s}$ supplies fine-grained textural priors (what).
+To effectively adapt to 3D medical data, we introduce customized operations (using MDiT3D-B as a representative example).
+Specifically, we apply a 3D patchify operator to project inputs into tokens $\mathbf{h}\in\mathbb{R}^{m\times T\times d}$ Mo et al.
+(2023), where $m$ is the number of modalities/slices and $d$ is the embedding dimension.
+Unlike standard 2D implementations Dosovitskiy et al.
+(2020), we incorporate 3D frequency-based sine-cosine positional embeddings (3D PE) to preserve precise spatial relationships.
 Alternating Blocks & Prompt Injection.
-Two distinct types of blocks are designed tailored for brain and cardiac tasks, respectively.
-For the brain MRI synthesis task, we adopt spatial and modal blocks: the former focuses on capturing 3D spatial information, while the latter models inter-modal relationships.
-Before each modal block, the tokens $\mathbf{h}$ are reshaped into $\mathbb{R}^{T\times m\times d}$ to facilitate inter-modal interaction, then reshaped back for subsequent spatial processing.
-These two types of blocks are stacked in an alternating fashion to form a total of $N$ blocks.
-The motivation behind this design lies in the need for the model to capture both inter- and intra-modal relationships to boost contextual understanding and anatomical consistency.
-Additionally, adaptive layer normalization is employed to inject prompts into the blocks.
-Notably, inspired by the intuition that conditioning signals should be subtle yet informative, prompt tokens are injected only into the modal blocks.
+We design two distinct alternating block architectures tailored to the specific characteristics of brain and cardiac MRIs.
+For the multi-modal brain task, we alternate between Spatial Blocks (to capture 3D spatial context) and Modal Blocks (to model inter-modal relationships).
+For the volumetric cardiac task, we alternate between Planar Blocks (for intra-slice features) and Spatial Blocks (for through-plane continuity).
+To compute attention along the appropriate axes, the latent tokens are dynamically reshaped before entering each block (e.g., aligning the modal dimension to facilitate inter-modal interaction, and then reshaped back for spatial processing).
+Furthermore, to maximize the efficacy of conditional guidance, prompts are injected via adaptive layer normalization (adaLN) Peebles & Xie (2023) exclusively into the blocks that explicitly model the task’s primary dependency.
+Specifically, prompt tokens are injected only into the Modal Blocks for brain MRI (addressing missing modalities) and the Spatial Blocks for cardiac MRI (addressing missing slices).
+This targeted injection avoids overwhelming the network and ensures the conditioning signals are both physically meaningful and highly informative.
 Joint Reconstruction & Synthesis.
-We concatenate the available and the noisy latent (corresponding to missing sections) as the input to MDiT3D.
-Moreover, a joint reconstruction & synthesis scheme is introduced, where only the missing sections undergo noise injection and denoising during the diffusion process, while the observed regions remain unchanged, serving as auxiliary context that provides rich semantic guidance for generation.
-Given that directly synthesizing the entire modality/volume is more informative and efficient than simply predicting zero noise for the available latent regions [29], we adopt x-prediction loss as objective function for MDiT3D, as below:
-$\small\mathcal{L}_{\text{diff}}=\mathbb{E}_{\mathbf{z}_{0},\epsilon,t}\left[\left\|\mathbf{z}_{0}-f_{\theta}(\mathbf{z}_{t},t,\mathbf{p})\right\|_{2}^{2}\right].$ (5)
-where $\mathbf{z}_{0}=\mathbf{z}^{s}\,\|\,\mathbf{z}^{c}$ denotes the concatenated clean latent tokens, $\mathbf{z}_{t}=\mathbf{z}^{s}\,\|\,\mathbf{z}_{t}^{c}$ is the noisy input at timestep $t$, $\mathbf{p}$ is the prompt tokens, and $f_{\theta}$ represents our MDiT3D model.
-4 Experiments and Results
-Table 2: Quantitative results for Cardiac MRI synthesis on UKBB dataset. The numbers of the second row denote the length of missing slices. The best performances are bolded.
-UKBB
+During diffusion, rather than filling missing modalities or slices with zeros or learned tokens, we only add noise to the missing sections, keeping the available latents unperturbed to provide rich contextual guidance.
+Following Meng et al.
+(2024), MDiT3D is optimized via an $\mathbf{x}_{0}$-prediction loss:
+$\mathcal{L}_{\text{diff}}=\mathbb{E}_{\mathbf{z}_{0},\epsilon,t}\left[\left\|\mathbf{z}_{0}-f_{\theta}(\mathbf{z}_{t},t,\mathbf{p})\right\|_{2}^{2}\right],$ (5)
+where $\mathbf{z}_{0}=\mathbf{z}^{s}\,\|\,\mathbf{z}^{c}$ is the clean target, $\mathbf{z}_{t}=\mathbf{z}^{s}\,\|\,\mathbf{z}_{t}^{c}$ is the partially noised input at timestep $t$, $\mathbf{p}$ represents the prompt tokens, and $f_{\theta}$ denotes MDiT3D.
+Table 3: Quantitative results for Cardiac MRI synthesis on the UKBB dataset. The numbers in the first row denote the length of missing slices.
 8 16 24
@@
 GAN-based Methods
-MMGAN [43] 25.81 $\pm$0.86 0.815 $\pm$0.012 16.68 $\pm$0.93 24.35 $\pm$0.84 0.793 $\pm$0.009 27.52 $\pm$0.90 23.06 $\pm$1.08 0.776 $\pm$0.018 45.88 $\pm$1.32
-MMT [24] 26.02 $\pm$0.82 0.824 $\pm$0.009 15.90 $\pm$0.88 24.73 $\pm$0.87 0.809 $\pm$0.010 23.75 $\pm$0.97 24.12 $\pm$1.06 0.794 $\pm$0.017 37.39 $\pm$1.27
-Hyper-GAE [53] 25.23 $\pm$0.89 0.810 $\pm$0.012 19.02 $\pm$0.96 23.66 $\pm$0.94 0.789 $\pm$0.013 31.84 $\pm$1.15 22.70 $\pm$1.14 0.771 $\pm$0.019 48.27 $\pm$1.39
+MMGAN Sharma & Hamarneh (2020) 25.81 $\pm$0.86 0.815 $\pm$0.012 16.68 392.48 24.35 $\pm$0.84 0.793 $\pm$0.009 27.52 604.37 23.06 $\pm$1.08 0.776 $\pm$0.018 45.88 767.92
+MMT Liu et al. (2023) 26.02 $\pm$0.82 0.824 $\pm$0.009 15.90 429.72 24.73 $\pm$0.87 0.809 $\pm$0.010 23.75 570.29 24.12 $\pm$1.06 0.794 $\pm$0.017 37.39 708.68
+Hyper-GAE Yang et al. (2023) 25.23 $\pm$0.89 0.810 $\pm$0.012 19.02 517.55 23.66 $\pm$0.94 0.789 $\pm$0.013 31.84 648.35 22.70 $\pm$1.14 0.771 $\pm$0.019 48.27 845.14
 Diffusion Model-based Methods
-LDM [39] 24.17 $\pm$0.96 0.795 $\pm$0.012 24.61 $\pm$0.94 23.04 $\pm$1.03 0.778 $\pm$0.010 44.93 $\pm$1.49 22.19 $\pm$1.21 0.761 $\pm$0.018 60.02 $\pm$1.68
-ControlNet [58] 24.62 $\pm$0.91 0.801 $\pm$0.012 22.47 $\pm$1.06 23.30 $\pm$0.98 0.784 $\pm$0.014 37.51 $\pm$1.55 22.46 $\pm$1.16 0.765 $\pm$0.020 54.93 $\pm$1.74
-M2DN [29] 25.48 $\pm$0.79 0.814 $\pm$0.011 17.72 $\pm$0.85 24.62 $\pm$0.87 0.803 $\pm$0.011 24.15 $\pm$1.06 24.03 $\pm$0.99 0.780 $\pm$0.021 40.08 $\pm$1.59
-CoPeDiT 26.42 $\pm$0.81 0.831 $\pm$0.013 15.53 $\pm$0.88 26.07 $\pm$0.74 0.826 $\pm$0.013 18.21 $\pm$0.90 25.39 $\pm$0.86 0.817 $\pm$0.016 25.84 $\pm$1.22
-Figure 4: Qualitative results on BraTS dataset.
-The results are depicted in the axial (top), sagittal (middle) and coronal (bottom) views of the 3D MRI volume.
-The visual results on IXI dataset are provided in the supplementary material.
-Figure 5: Qualitative results of comparison on UKBB dataset. The top and bottom results correspond to the first and last missing slices within a given volume, respectively.
-Table 3: Ablation study on the contribution of each pretext task to CoPeVAE’s reconstruction capacity.
-‘Cardiac’ refers to the aggregated cardiac MRI dataset consisting of four sources, which are used for training our CoPeVAE model.
+LDM Rombach et al. (2022) 24.17 $\pm$0.96 0.795 $\pm$0.012 24.61 704.91 23.04 $\pm$1.03 0.778 $\pm$0.010 44.93 830.73 22.19 $\pm$1.21 0.761 $\pm$0.018 60.02 910.15
+ControlNet Zhang et al. (2023) 24.62 $\pm$0.91 0.801 $\pm$0.012 22.47 698.63 23.30 $\pm$0.98 0.784 $\pm$0.014 37.51 821.40 22.46 $\pm$1.16 0.765 $\pm$0.020 54.93 922.06
+M2DN Meng et al. (2024) 25.48 $\pm$0.79 0.814 $\pm$0.011 17.72 453.03 24.62 $\pm$0.87 0.803 $\pm$0.011 24.15 597.78 24.03 $\pm$0.99 0.780 $\pm$0.021 40.08 823.70
+DiffM4RI Ye et al. (2026) 25.19 $\pm$0.89 0.808 $\pm$0.014 19.16 495.23 24.78 $\pm$0.80 0.807 $\pm$0.015 22.96 548.44 24.27 $\pm$1.06 0.791 $\pm$0.018 35.27 674.12
+CoPeDiT 26.42 $\pm$0.81 0.831 $\pm$0.013 15.53 318.62 26.07 $\pm$0.74 0.826 $\pm$0.013 18.21 382.04 25.39 $\pm$0.86 0.817 $\pm$0.016 25.84 490.57
+4 Experiments and Results
+4.1 Experimental Setup
+Datasets.
+(i) Brain MRI Datasets.
+We evaluate the effectiveness of CoPeDiT on two public brain MRI datasets: BraTS 2021 Baid et al.
+(2021) and IXI Brain Development Project (2025).
+The BraTS 2021 dataset includes 1251 subjects with multi-modal MRI scans across four modalities: T1, T1ce, T2, and FLAIR.
+The IXI dataset contains 577 subjects with three MRI modalities: T1, T2, and PD.
+(ii) Cardiac MRI Datasets.
+Missing slice synthesis experiments are conducted on four cardiac MRI datasets: UK Biobank (UKBB) Petersen et al.
+(2016), MESA Zhang et al.
+(2018), ACDC Bernard et al.
+(2018), and MSCMR Zhuang et al.
+(2022).
+The model is trained on the combined dataset including all four sources with 32,248 MRI volumes in total, while performance comparisons are conducted on the UKBB dataset.
+We randomly select 80% of the data for training and use the remaining 20% as the test set.
+Please refer to Appendix A for more details.
+Implementation Details.
+The compression rate of CoPeVAE is set to $(8,8,8)$.
+For hyperparameters, the dimension of each prompt token is set to 512, $\tau$ and $\lambda$ are set to 0.2 and 1e-2, respectively.
+The model is trained with a global batch size of 8 for CoPeVAE-B and 64 for CoPeVAE-C, using a learning rate of 1e-4.
+Regarding MDiT3D, we set the number of time steps to 500 with linearly scaled noise scheduling.
+The model is trained for 100k steps with a global batch size of 32 for MDiT3D-B and 64 for MDiT3D-C, with learning rate of 5e-5.
+All training is conducted on four NVIDIA A100 GPUs.
+More details are provided in Appendix B.
+Figure 4: Qualitative results on the BraTS dataset. The visual results on the IXI dataset are provided in Appendix C.1.
+Figure 5: Qualitative results on the UKBB dataset. The top and bottom results correspond to the first and last missing slices within a given volume, respectively.
+Table 4: Ablation study on the contribution of completeness-aware prompts.
+BraTS UKBB
+1 2 3 8 24
+[…]
+w/o $\mathbf{p}^{d}$ 27.35 0.833 16.04 335.08 27.02 0.824 17.30 475.32 26.73 0.814 19.62 556.79 25.82 0.819 16.34 427.06 24.53 0.803 33.80 610.46
+w/o $\mathbf{p}^{p}$ 26.92 0.829 18.46 372.36 26.43 0.819 20.23 460.45 26.06 0.810 25.13 572.03 25.17 0.812 20.08 534.80 24.39 0.798 36.54 697.51
+w/o $\mathbf{p}^{s}$ 27.56 0.835 15.26 341.70 27.22 0.827 16.37 395.43 26.90 0.815 17.58 468.23 26.27 0.828 16.79 388.46 25.08 0.813 29.83 578.13
+w/o Prompts 25.92 0.823 25.69 418.12 25.06 0.807 32.17 556.92 24.83 0.802 37.97 748.25 24.70 0.797 23.29 721.35 23.56 0.778 42.17 830.72
+w/ Mask Codes 27.18 0.831 17.42 356.09 26.82 0.816 20.06 439.13 26.18 0.809 24.59 543.82 26.15 0.823 18.15 409.47 24.65 0.802 35.86 612.73
+CoPeDiT 28.26 0.842 12.67 254.71 28.13 0.831 13.25 287.58 27.91 0.822 14.89 323.19 26.42 0.831 15.53 318.62 25.39 0.817 25.84 490.57
+Table 5: Quantitative results by incorporating our learned prompt tokens into baselines instead of mask codes.
+[…]
+MMT Liu et al. (2023) 25.19 0.824 24.53 527.06
++ Prompts (ours) 25.68 (+0.49) 0.826 (+0.002) 22.07 (–2.46) 416.34 (–110.72)
+Hyper-GAE Yang et al. (2023) 24.65 0.813 28.97 609.65
++ Prompts (ours) 25.26 (+0.61) 0.822 (+0.009) 24.26 (–4.71) 523.75 (–85.90)
+M2DN Meng et al. (2024) 26.45 0.830 21.29 376.53
++ Prompts (ours) 27.23 (+0.78) 0.837 (+0.007) 17.06 (–4.23) 308.68 (–67.85)
+Table 6: Ablation study of each pretext task’s contribution to CoPeVAE’s reconstruction capacity. ’Cardiac’ refers to the combined dataset.
 BraTS IXI Cardiac
@@
 CoPeVAE 35.05 0.935 31.28 0.921 33.34 0.931
-Table 4: Ablation study on the contribution of completeness-aware prompt tokens.
+4.2 Performance Comparison
+We compare our method against seven SOTA baselines, including three GAN-based approaches Sharma & Hamarneh (2020); Liu et al.
+(2023); Yang et al.
+(2023) and four diffusion-based models Rombach et al.
+(2022); Zhang et al.
+(2023); Meng et al.
+(2024); Ye et al.
+(2026), all reimplemented under identical settings for fair comparison.
+Performance is evaluated using Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM), Fréchet Inception Distance (FID) Heusel et al.
+(2017), and Fréchet Video Distance (FVD) Unterthiner et al.
+(2018) for assessing 3D spatial consistency.
+Quantitative Results.
+Quantitative results on the three datasets are presented in Tables 12 and 3, respectively.
+CoPeDiT consistently outperforms all baselines across all missing configurations in both synthesis tasks.
+Notably, these performance gains widen in scenarios with a higher number of missing modalities or slices, such as maintaining a high 27.91 PSNR even with three missing modalities.
+This highlights the robustness of our completeness-aware prompts in complex cases.
+Moreover, CoPeDiT achieves substantially lower FID and FVD scores (e.g., an FVD of 490.57 for 24 missing cardiac slices ).
+This indicates that the generated MRIs are not only anatomically coherent and texture-preserving, but exhibit superior 3D spatial consistency and structural continuity, enhancing perceptual realism and diagnostic plausibility.
+Qualitative Results.
+As depicted in Figs.
+4 and 5, our model generates synthetic MRIs that exhibit the highest visual similarity to the ground truth images, particularly in accurately capturing tumor regions.
+Our CoPeDiT excels at preserving subtle textural details and modeling complex anatomical structures within brain tissues, justifying our motivation that incorporating completeness perception leads to improved anatomical consistency and realism.
+Table 7: Ablation study on the choice of prompt injection positions within the MDiT3D blocks.
+The labels denote the target blocks for brain and cardiac tasks, respectively (e.g., "Modal / Spatial" means injecting prompts exclusively into the modal blocks for brain MRIs and spatial blocks for cardiac MRIs).
+"Both" indicates injection into both types of blocks for the respective tasks.
 BraTS UKBB
@@
 […]
-w/o $\mathbf{p}^{d}$ 27.35 0.833 16.04 27.02 0.824 17.30 26.73 0.814 19.62 25.82 0.819 16.34 24.53 0.803 33.80
-w/o $\mathbf{p}^{p}$ 26.92 0.829 18.46 26.43 0.819 20.23 26.06 0.810 25.13 25.17 0.812 20.08 24.39 0.798 36.54
-w/o $\mathbf{p}^{s}$ 27.56 0.835 15.26 27.22 0.827 16.37 26.90 0.815 17.58 26.27 0.828 16.79 25.08 0.813 29.83
-w/o Prompts 25.92 0.823 25.69 25.06 0.807 32.17 24.83 0.802 37.97 24.70 0.797 23.29 23.56 0.778 42.17
-w/ Mask Codes 27.18 0.831 17.42 26.82 0.816 20.06 26.18 0.809 24.59 26.15 0.823 18.15 24.65 0.802 35.86
-CoPeDiT 28.26 0.842 12.67 28.13 0.831 13.25 27.91 0.822 14.89 26.42 0.831 15.53 25.39 0.817 25.84
-Table 5: Quantitative results by incorporating our learned prompt tokens into baselines instead of mask codes.
-BraTS
-[…]
-MMT [24] 25.19 0.824 24.53
-+ Prompts (ours) 25.68 (+0.49) 0.826 (+0.002) 22.07 (–2.46)
-Hyper-GAE [53] 24.65 0.813 28.97
-+ Prompts (ours) 25.26 (+0.61) 0.822 (+0.009) 24.26 (–4.71)
-M2DN [29] 26.45 0.830 21.29
-+ Prompts (ours) 27.23 (+0.78) 0.837 (+0.007) 17.06 (–4.23)
-4.1 Experimental Setup
-Datasets.
-(i) Brain MRI Datasets.
-We evaluate the effectiveness of CoPeDiT on two public brain MRI datasets: BraTS 2021 [3] and IXI [5].
-The BraTS 2021 dataset includes 1251 subjects with multi-modal MRI scans across four modalities: T1, T1ce, T2, and FLAIR.
-The IXI dataset contains 577 subjects with three MRI modalities: T1, T2, and PD.
-(ii) Cardiac MRI Datasets.
-Missing slice synthesis experiments are conducted on four cardiac MRI datasets: UK Biobank (UKBB) [35], MESA [57], ACDC [4], and MSCMR [59].
-The model is trained on the combined dataset including all four sources with 32,248 MRI volumes in total, while performance comparisons are conducted on the UKBB dataset.
-We randomly select 80% of the data for training and use the remaining 20% as the test set.
-Please refer to the supplementary material for more dataset details.
-Implementation Details.
-The compression rate of CoPeVAE is set to $(8,8,8)$.
-For hyperparameters, the dimension of each prompt token is set to 512, $\tau$ and $\lambda$ are set to 0.2 and 1e-2, respectively.
-The model is trained with a global batch size of 8 for CoPeVAE-B and 64 for CoPeVAE-C, using a learning rate of 1e-4.
-Regarding MDiT3D, we set the number of time steps to 500 with linearly scaled noise scheduling.
-The model is trained for 100k steps with a global batch size of 32 and learning rate of 5e-5.
-All training is conducted on four NVIDIA A100 GPUs.
-More details are provided in the supplementary material.
-4.2 Performance Comparison
-Quantitative Results.
-We compare our method against six recent SOTA models for 3D brain and cardiac MRI synthesis, including three GAN-based approaches [43, 24, 53] and three advanced diffusion model-based methods [39, 58, 29].
-We report results by reimplementing all baselines from scratch under identical experimental settings to ensure fair comparisons.
-Performance is evaluated using three metrics: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM), and Fréchet Inception Distance (FID) [15] that uses Inception-V3 [46] as the feature extractor.
-Quantitative results on the three datasets are presented in Tables 1 and  2, where the second row indicates the number of missing modalities or slices $c$, respectively.
-The results show that CoPeDiT consistently outperforms all baselines across all missing configurations in both synthesis tasks, showcasing superior generalizability.
-Notably, the performance gains are more pronounced in scenarios with a larger number of missing modalities/slices, highlighting the robustness of our completeness-aware prompts in handling complex missing cases.
-Besides, CoPeDiT achieves significantly lower FID, indicating that the generated MRIs are more anatomically coherent and texture-preserving, with feature distributions that are closer to real clinical data, thus enhancing perceptual realism and diagnostic plausibility.
-Qualitative Results.
-As depicted in Fig.
-4 and 5, our model generates synthetic MRIs that exhibit the highest visual similarity to the ground truth images, particularly in accurately capturing tumor regions.
-Our CoPeDiT excels at preserving subtle textural details and modeling complex anatomical structures within brain tissues, justifying our motivation that incorporating completeness perception leads to improved anatomical consistency and realism.
-Table 6: Ablation study on the choice of prompt injection positions within the MDiT3D blocks.
-‘Spatial / Planar’ indicates that prompt tokens are injected into the spatial and planar blocks for brain and cardiac tasks, respectively.
-‘Both’ indicates that prompt tokens are injected into both the spatial and modal blocks for brain tasks, and into both the planar and spatial blocks for cardiac tasks.
-‘Modal / Spatial’ indicates that prompt tokens are injected into the modal and spatial blocks, respectively.
-BraTS UKBB
-1 2 3 8 24
-[…]
-Spatial / Planar 27.47 0.834 15.41 27.20 0.825 16.83 27.03 0.817 17.45 26.29 0.827 16.82 25.22 0.814 29.24
-Both 27.76 0.836 13.94 27.54 0.827 15.48 27.29 0.818 16.77 26.33 0.827 16.39 25.28 0.815 28.45
-Modal / Spatial 28.26 0.842 12.67 28.13 0.831 13.25 27.91 0.822 14.89 26.42 0.831 15.53 25.39 0.817 25.84
+Spatial / Planar 27.47 0.834 15.41 317.62 27.20 0.825 16.83 484.39 27.03 0.817 17.45 592.17 26.29 0.827 16.82 405.72 25.22 0.814 29.24 604.32
+Both 27.76 0.836 13.94 294.37 27.54 0.827 15.48 367.20 27.29 0.818 16.77 434.83 26.33 0.827 16.39 414.37 25.28 0.815 28.45 587.13
+Modal / Spatial 28.26 0.842 12.67 254.71 28.13 0.831 13.25 287.58 27.91 0.822 14.89 323.19 26.42 0.831 15.53 318.62 25.39 0.817 25.84 490.57
 Table 8: Ablation study on the contribution of 3D MRI Diffusion Transformer.
@@
 […]
-UNet [39] 25.61 0.825 23.56 25.20 0.816 28.37 24.78 0.805 36.70 24.46 0.800 22.34 23.01 0.778 41.59
-DiT [34] 26.78 0.835 19.72 26.14 0.823 21.84 25.74 0.808 24.47 25.89 0.825 16.28 24.78 0.799 36.10
-DiT-3D [30] 26.84 0.835 19.23 26.23 0.822 21.95 25.90 0.810 26.08 26.08 0.825 16.37 24.83 0.806 33.73
-MDiT3D 28.26 0.842 12.67 28.13 0.831 13.25 27.91 0.822 14.89 26.42 0.831 15.53 25.39 0.817 25.84
+UNet Rombach et al. (2022) 25.61 0.825 23.56 433.55 25.20 0.816 28.37 545.89 24.78 0.805 36.70 707.83 24.46 0.800 22.34 689.53 23.01 0.778 41.59 871.64
+DiT Peebles & Xie (2023) 26.78 0.835 19.72 392.09 26.14 0.823 21.84 498.13 25.74 0.808 24.47 611.26 25.89 0.825 16.28 403.83 24.78 0.799 36.10 592.72
+DiT-3D Mo et al. (2023) 26.84 0.835 19.23 359.83 26.23 0.822 21.95 487.12 25.90 0.810 26.08 593.42 26.08 0.825 16.37 397.71 24.83 0.806 33.73 610.89
+MDiT3D 28.26 0.842 12.67 254.71 28.13 0.831 13.25 287.58 27.91 0.822 14.89 323.19 26.42 0.831 15.53 318.62 25.39 0.817 25.84 490.57
 4.3 Ablation Study
+Effect of Prompt Tokens.
+Table 4 compares learned prompt tokens against one-hot mask codes.
+Under identical injection strategies, all individual prompts and their combinations consistently outperform binary masks, strongly validating our completeness perception approach.
+Specifically, the positioning prompt ($p^{p}$) proves most effective; its removal triggers the sharpest performance drop, likely because explicitly locating missing sections heightens sensitivity to subtle structural variations.
+To further explore the potential of our learned prompts, we incorporate them into prior baselines as replacements for mask codes.
+As illustrated in Table 6, this substitution consistently boosts their performance, such as yielding a 0.78 dB PSNR gain for M2DN, demonstrating the strong plug-and-play utility and generalizability of our method.
 Effect of Pretext Tasks.
@@
 This improvement can be attributed to the fact that pretext tasks promote the tokenizer to capture anatomical structure variances in both coarse and fine-grained manner, learning highly discriminative features.
-Effect of Prompt Tokens.
-Table 4 presents the synthesis performance associated with each learned prompt token and mask code.
-The mask codes are generated as one-hot vectors whose dimension corresponds to the number of modalities or slices, indicating the corresponding missing positions.
-Notably, we employ an identical injection strategy for all prompts and masks to ensure a fair comparison.
-As shown, all individual prompt tokens and their combinations improve model performance, surpassing guidance based on binary mask codes.
-These findings strongly highlight the effectiveness and rationale of our completeness perception scheme.
-The incompleteness positioning task yields the most effective prompts, likely due to its explicit emphasis on positioning missing sections along the modality/slice dimension, thus increasing the sensitivity to subtle structural variations.
-To further explore the potential of our learned prompts, we incorporate them into prior baselines as replacements for mask codes.
-As illustrated in Table 13, this substitution consistently improves their performance, demonstrating the utility and generalizability of our completeness-aware prompts, which can serve as a flexible paradigm for improving emerging MRI synthesis methods.
 Choice of Prompt Injection Position.
-This experiment aims to justify our choice of injecting prompts exclusively into the modal and spatial blocks for brain and cardiac tasks, respectively.
-As illustrated in Table 14, injecting prompts into the modal/spatial block yields the best overall performance, whereas injecting into spatial/planar blocks or both blocks results in inferior outcomes.
+We aim to justify our choice of injecting prompts exclusively into the modal and spatial blocks for brain and cardiac tasks, respectively.
+As shown in Table 7, injecting prompts into the modal/spatial block yields the best performance, whereas injecting into spatial/planar blocks or both blocks results in inferior outcomes.
 This can be attributed to the fact that aligning the prompt with the block that explicitly models the task’s primary dependency, namely modality fusion for brain and through-plane continuity for cardiac, maximizes the efficacy of conditional signals.
 Impact of 3D MRI Diffusion Transformer.
-The quantitative evaluation of MDiT3D and existing diffusion networks [39, 34, 30] is presented in Table 7.
-The results illustrate that MDiT3D consistently achieves superior performance, confirming that the specialized architectural designs effectively enhance the quality of 3D MRI synthesis.
+The evaluation of MDiT3D and existing diffusers Rombach et al.
+(2022); Peebles & Xie (2023); Mo et al.
+(2023) is presented in Table 8.
+MDiT3D achieves superior performance, notably outperforming vanilla 3D DiTs by over 1.4 dB in PSNR on the BraTS dataset.
+This confirms that our task-driven design, which couples alternating blocks with targeted prompt injection to explicitly model inter-modal relationships and through-plane continuity, effectively improves synthesis.
+Table 9: Results of tumor segmentation experiments on the BraTS dataset.
+[…]
+WT TC ET AVG
+Missing 86.08 84.67 81.59 84.11
+GAN-based Methods
+MMGAN Sharma & Hamarneh (2020) 89.35 88.14 87.73 88.41
+MMT Liu et al. (2023) 90.43 88.37 86.92 88.57
+Hyper-GAE Yang et al. (2023) 88.72 86.54 85.37 86.88
+Diffusion Model-based Methods
+LDM Rombach et al. (2022) 87.86 85.91 84.19 85.99
+ControlNet Zhang et al. (2023) 88.27 87.05 85.23 86.85
+M2DN Meng et al. (2024) 91.28 90.09 88.20 89.86
+DiffM4RI Ye et al. (2026) 90.04 89.23 87.68 88.72
+CoPeDiT 91.35 90.41 88.94 90.23
+Figure 6: Visualization of salient regions on the BraTS dataset.
+4.4 Tumor Segmentation
+To evaluate clinical utility, we conduct downstream tumor segmentation on the BraTS dataset.
+Following Liu et al.
+(2023); Meng et al.
+(2024), we train a multi-modal U-Net Ronneberger et al.
+(2015) using three available modalities alongside one modality synthesized by CoPeDiT or baselines.
+We also evaluate a "Missing" baseline trained solely on the three available modalities.
+As shown in Table 9, all synthesis methods improve upon the "Missing" baseline in terms of average Dice scores for whole tumor (WT), tumor core (TC), and enhancing tumor (ET).
+Notably, CoPeDiT consistently outperforms competing methods across all subregions, achieving the highest average Dice of 90.23%.
+These results confirm that our synthesized MRIs provide highly informative and clinically valuable inputs for downstream tasks.
+4.5 Computational Efficiency Analysis
+Table 10: The computational cost and wall-clock time comparison of CoPeVAE on the BraTS dataset.
+Param. (M) Flops (G) Epoch Time (s) Total Time (GPU hours) BraTS (Avg)
+[…]
+w/o Task 1 136.77 21167.61 331.3 561.2 27.03 455.73
+w/o Task 2 136.55 21171.83 332.9 563.9 26.47 468.13
+w/o Task 3 136.74 21171.83 330.7 560.1 27.23 401.79
+CoPeVAE 137.63 21184.57 336.7 570.3 28.10 288.49
+Table 11: The inference cost comparison of CoPeDiT on the BraTS dataset.
+Inference Latency (s) VRAM (G)
+LDM Rombach et al. (2022) 1.38 0.429
+M2DN Meng et al. (2024) 4.67 0.819
+DiffM4RI Ye et al. (2026) 2.16 0.584
+DiT-3D Mo et al. (2023) 1.65 0.614
+CoPeDiT 1.73 0.626
+Training Compute and Wall-Clock Cost.
+As shown in Table 4.5, our CoPeVAE adds minimal overhead in parameters, FLOPs, and training time compared to variants.
+This modest cost yields superior synthesis quality and spatial consistency, proving a highly efficient complexity-gain trade-off for our pretext tasks.
+Inference Cost.
+To assess practical applicability, Table 4.5 reports end-to-end latency and VRAM usage per volume under standardized settings (batch size = 1, single A100 GPU, DDIM Song et al.
+(2021) 200 steps, mixed precision).
+CoPeDiT remains efficient, with resource demands comparable to standard models and markedly lower than heavier baselines, supporting its suitability for clinical deployment.
+Figure 7: Qualitative analysis of the learned prompts on the BraTS dataset.
+(a) t-SNE projection of the count-focused prompt $\mathbf{p}^{d}$, showing clear clustering by the number of missing modalities.
+(b) t-SNE projection of the identity-focused prompt $\mathbf{p}^{p}$, revealing distinct separation by missing modality types.
+(c) Visualization of modal block attention maps under a missing configuration of [1, 0, 0, 0].
 4.6 Visualization Analysis
-Figure 6: Visualization of salient regions on BraTS dataset.
-Figure 7: t-SNE visualization on the BraTS dataset of (a) The count-focused prompt $\mathbf{p}^{d}$ colored by missing-number classes; and (b) The identity-focused prompt $\mathbf{p}^{p}$ colored by missing modality.
-Figure 8: Visualization of attention maps from individual modal blocks on BraTS dataset under a missing configuration of [1, 0, 0, 0], where the first modality is absent.
 Salient Regions.
-To better understand the learning procedure of pretext tasks, we visualize their activation maps using GradCAM [42].
+To understand the learning procedure of pretext tasks, we visualize their activation maps using GradCAM Selvaraju et al.
+(2017).
 As shown in Fig.
-6, the four examples reveal strong correlations between salient regions and modality-discriminative features.
-Specifically, these patterns mirror each task’s objective: Task 1 and Task 3 must assess global consistency, so they rely on coarse anatomical layout (gray and white matter) that summarize the volume.
-In contrast, Task 2 needs to pinpoint where/which is missing, so it keys on high-frequency, modality-specific cues (tumors, lesions, and white matter hyperintensities in FLAIR) to localize.
-The results emphasize that our pretext tasks effectively capture modality-specific properties, understanding the MRI context and relationships.
+6, the examples reveal strong correlations between salient regions and modality-discriminative features.
+These patterns mirror each task’s objective: Task 1 and Task 3 must assess global consistency, so they rely on coarse anatomical layout (gray and white matter) that summarize the volume.
+In contrast, Task 2 needs to pinpoint where/which is missing, so it keys on high-frequency, modality-specific cues (tumors, lesions, and white matter hyperintensities in FLAIR).
+The results emphasize that our pretext tasks capture modality-specific properties and improve the model’s understanding of MRI context and inter-modality relationships.
 Prompt Distribution.
 We implement the t-SNE visualization of learned prompts on BraTS and color them by ground-truth incompleteness labels.
-As depicted in Fig.
-7, the count-focused prompt $\mathbf{p}^{d}$ forms compact, well-separated clusters aligned with the missing number classes corresponding to each missing state (1/2/3 absent modalities), while the identity-focused prompt $\mathbf{p}^{p}$ produces modality-specific clusters (T1, T1ce, T2, FLAIR) with clear boundaries, highlighting that the two prompts capture complementary aspects of incompleteness.
+As depicted in Figs.
+7 and 7, the count-focused prompt $\mathbf{p}^{d}$ forms compact, well-separated clusters aligned with the missing number classes corresponding to each missing state (1/2/3 absent modalities).
+Meanwhile, the identity-focused prompt $\mathbf{p}^{p}$ produces modality-specific clusters (T1, T1ce, T2, FLAIR) with distinct boundaries.
+This clear separation visually confirms that our pretext tasks effectively drive the tokenizer to learn highly discriminative features.
+Together, they demonstrate that our learned prompt tokens capture explicitly decoupled and complementary semantic priors.
+By avoiding feature entanglement, these prompts successfully provide the network with nuanced, multi-granular guidance for reliable synthesis.
 Attention Maps.
-We investigate the contribution of prompts by visualizing the attention maps of each modal block.
+We visualize the attention maps of each modal block under a specific missing configuration (e.g., [1, 0, 0, 0]).
 As shown in Fig.
-8, our prompts guide the attention mechanism to progressively focus on the actual missing elements (the first row and column) with the increase of block layers.
-In comparison, mask codes lack sufficient informativeness and fail to provide nuanced guidance that aligns with the true missing state.
+7, our learned prompts actively guide the attention mechanism to progressively focus on the actual missing elements (the first row and column) as the block depth increases.
+This concentrated focus facilitates precise, layer-wise feature aggregation for the absent modality.
+In comparison, explicit mask codes lack sufficient informativeness, yielding diffuse attention weights that fail to provide the nuanced, dynamic guidance necessary to align with the true missing state.
+Figure 8: Quantitative analysis of completeness-aware prompts on the BraTS dataset.
+(a) Sensitivity of CoPeDiT to increasing prompt corruption rates.
+Synthesis quality degrades monotonically, particularly when perturbing the position prompt $\mathbf{p}^{p}$.
+(b) Prompt intervention matrix illustrating the PSNR drop under structured perturbations.
+Misleading signals (e.g., random, type-swap) and the simultaneous perturbation of all prompts cause the most severe performance degradation.
+4.7 Analyzing Completeness-Aware Prompts
+To quantitatively validate the effectiveness and rationale of our learned prompts, we conduct comprehensive sensitivity and intervention analyses.
+Prompt Accuracy Sensitivity.
+To assess CoPeDiT’s reliance on prompt accuracy, we randomly replace the degree ($\mathbf{p}^{d}$), position ($\mathbf{p}^{p}$), semantic ($\mathbf{p}^{s}$), or "All" prompts with incorrect ones for an $r$-fraction ($r\in[0,1.0]$) of BraTS validation samples.
+Fig.
+8 demonstrates that synthesis quality degrades consistently as the corruption rate $r$ increases.
+Corrupting all prompts simultaneously yields the worst outcomes, confirming their complementary and essential roles.
+Prompt Intervention.
+We further apply structured interventions by replacing prompt components with zeroed, random, shuffled, or type-swapped variants.
+As Fig.
+8 illustrates, misleading signals (random/type-swapped) degrade performance more severely than missing or mildly perturbed guidance (zeroed/shuffled).
+Most notably, across both sensitivity and intervention experiments, perturbing $\mathbf{p}^{p}$ consistently triggers the sharpest performance drops.
+This combined evidence robustly demonstrates that precise missing region localization ($\mathbf{p}^{p}$) is the most crucial prompt for guiding high-fidelity 3D MRI synthesis.
 5 Conclusion
 This work presents CoPeDiT, a unified model for 3D brain and cardiac MRI synthesis that explores completeness perception.
-We demonstrate that enabling the model to autonomously infer missing state, rather than relying on externally pre-defined masks, can provide more discriminative and informative guidance.
-To this end, we equip our tokenizer with completeness perception capability through elaborated pretext tasks.
+We demonstrate that enabling the model to autonomously infer the missing state, rather than relying on externally pre-defined masks, can provide more discriminative and informative guidance.
+To this end, we equip our tokenizer with completeness perception capability through carefully designed pretext tasks.
 MDiT3D is then developed to utilize the learned prompt tokens as guidance for 3D MRI generation.
-Experimental evaluation validates the superiority of CoPeDiT in terms of accuracy, generalizability, robustness, and adaptability across diverse missing scenarios.
-Furthermore, CoPeDiT enables mask-free synthesis through perceiving completeness state, enhancing flexibility and applicability in practical clinical settings.
-Supplementary Material
+Extensive evaluations validate CoPeDiT’s remarkable accuracy and robustness across diverse scenarios, highlighting its potential for practical clinical deployment.
+Limitations.
+While highly effective, CoPeDiT requires a fixed number of modalities during training and may lose some fine high-frequency details due to latent space compression.
+Future work will focus on modality-agnostic tokenizers and exploring pixel-space diffusion refinement.
+Appendix
 Appendix A Datasets
@@
 Brain MRI
-BraTS [3] T1, T1ce, T2, FLAIR 1251 1000 251
-IXI [5] T1, T2, PD 577 462 115
+BraTS Baid et al. (2021) T1, T1ce, T2, FLAIR 1251 1000 251
+IXI Brain Development Project (2025) T1, T2, PD 577 462 115
 Cardiac MRI
-UKBB [35] - 31350 25080 6270
-MESA [57] - 298 238 60
-ACDC [4] - 300 240 60
-MSCMR [59] - 300 240 60
+UKBB Petersen et al. (2016) - 31350 25080 6270
+MESA Zhang et al. (2018) - 298 238 60
+ACDC Bernard et al. (2018) - 300 240 60
+MSCMR Zhuang et al. (2022) - 300 240 60
 Combined - 32248 25798 6450
@@
 Brain MRI data.
-Following [24, 29], we use 90 and 80 middle axial slices for BraTS and IXI datasets, respectively.
+Following Liu et al.
+(2023); Meng et al.
+(2024), we use 90 and 80 middle axial slices for BraTS and IXI datasets, respectively.
 These slices are further cropped to a size of $192\times 192$ from the central region.
@@
 The original binary masks that were used to generate synthetic missingness are not provided to the diffusion model.
-In this sense, our synthesis process is mask-free: CoPeDiT no longer depends on hand-crafted or externally supplied mask codes at the generation stage, but instead relies entirely on the learned completeness prompts produced by the frozen CoPeVAE.
+In this sense, CoPeDiT no longer depends on hand-crafted or externally supplied mask codes at the generation stage, but instead relies entirely on the learned completeness prompts produced by the frozen CoPeVAE.
 B.3 Hyperparameter Setups
@@
 The detailed hyperparameter setup of CoPeVAE is provided in Table 14.
-Built upon VQVAE [47] and VQGAN [10], our model employs a codebook containing 8192 codes with the latent dimensionality of 8.
+Built upon VQVAE van den Oord et al.
+(2017) and VQGAN Esser et al.
+(2021), our model employs a codebook containing 8192 codes with the latent dimensionality of 8.
 The values of $\tau$ and $\lambda$ are empirically set, as they have only a slight impact on model performance.
@@
 The training steps of CoPeVAE-B and CoPeVAE-C are 400k and 100k, respectively.
-Figure 9: Qualitative results of comparison on IXI dataset. The results are depicted in the axial (top), sagittal (middle) and coronal (bottom) views of the 3D MRI volume.
 Table 14: Hyperparameter setup of CoPeVAE.
@@
 LR schedule Warmup cosine Warmup cosine
-Training steps 400k 200k
+Training steps 400k 100k
 Table 15: Hyperparameter setup of MDiT3D.
@@
 Interpolants
-Training objective x-prediction x-prediction
+Training objective $\mathbf{x}_{0}$-prediction $\mathbf{x}_{0}$-prediction
 Noise schedule scaled-linear scaled-linear
@@
 For MDiT3D, we carefully design the hyperparameters to balance dataset size and model capacity, as summarized in Table 15.
-Following DiT [34], we report the experimental results using the exponential moving average (EMA) with a decay rate of 0.9999.
+Following DiT Peebles & Xie (2023), we report the experimental results using the exponential moving average (EMA) with a decay rate of 0.9999.
 During training, we set the timestep to 500 with linearly scaled noise levels ranging from 0.0015 to 0.0195.
 MDiT3D is trained for 100k iterations using the AdamW optimizer and a warmup cosine learning rate schedule.
-During inference, the DDIM sampler is applied with sample steps of 200 and 250 for MDiT3D-B and MDiT3D-C, respectively.
+During inference, the DDIM sampler Song et al.
+(2021) is applied with sample steps of 200 and 250 for MDiT3D-B and MDiT3D-C, respectively.
 In addition, we use mixed-precision training with a gradient clipping to accelerate training and save computational resources throughout all two-stage experiments.
 Appendix C Additional Experimental Results
+Figure 9: Qualitative results on the IXI dataset.
 C.1 Qualitative Results
@@
 As shown, we can also observe that our CoPeDiT outperforms other baselines in preserving the intricate structures and texture information in the synthesized MRIs.
-Figure 10: Sensitivity of CoPeDiT to prompt corruption on BraTS dataset.
-Table 12: Experimental results of tumor segmentation experiments on BraTS dataset.
-[…]
-WT TC ET AVG
-Missing 86.08 84.67 81.59 84.11
-GAN-based Methods
-MMGAN [43] 89.35 88.14 87.73 88.41
-MMT [24] 90.43 88.37 86.92 88.57
-Hyper-GAE [53] 88.72 86.54 85.37 86.88
-Diffusion Model-based Methods
-LDM [39] 87.86 85.91 84.19 85.99
-ControlNet [58] 88.27 87.05 85.23 86.85
-M2DN [29] 91.28 90.09 88.20 89.86
-CoPeDiT 91.35 90.41 88.94 90.23
-C.2 Tumor Segmentation
-In this section, we evaluate the clinical utility of our synthesized MRIs through downstream tumor segmentation experiments on the BraTS dataset.
-Specifically, following [24, 29], we train a multi-modal U-Net [40] on the four MRI modalities, where one modality is withheld and replaced by the synthesized image generated by our CoPeDiT or competing baselines.
-Additionally, we implement a U-Net that takes only the three available modalities as input for comparison.
-We report the average Dice scores for whole tumor (WT), tumor core (TC), and enhancing tumor (ET), as summarized in Table 12.
-As shown, all methods improve the segmentation accuracy over the ’Missing’ baseline that only uses the three available modalities, confirming the benefit of recovering the missing contrast.
-Among them, CoPeDiT achieves the best performance on all three tumor subregions, leading to the highest average Dice of 90.23%.
-These results demonstrate that our synthesized MRIs are not only visually plausible but also provide more informative inputs for downstream clinical tasks such as tumor segmentation.
 Table 16: Ablation study on the contribution of completeness perception prompt tokens on the IXI dataset.
-IXI
 1 2
 […]
-w/o $\mathbf{p}^{d}$ 23.56 0.720 35.93 23.04 0.708 48.71
-w/o $\mathbf{p}^{p}$ 22.81 0.697 50.70 21.99 0.692 63.16
-w/o $\mathbf{p}^{s}$ 24.17 0.724 31.89 23.56 0.710 39.87
-w/o Prompts 22.48 0.696 57.64 21.67 0.683 81.49
-w/ Mask Codes 23.35 0.711 44.03 22.84 0.702 56.72
-CoPeDiT 24.34 0.732 25.84 23.92 0.721 32.53
+w/o $\mathbf{p}^{d}$ 23.56 0.720 35.93 724.18 23.04 0.708 48.71 917.76
+w/o $\mathbf{p}^{p}$ 22.81 0.697 50.70 1267.70 21.99 0.692 63.16 1281.31
+w/o $\mathbf{p}^{s}$ 24.17 0.724 31.89 685.13 23.56 0.710 39.87 876.94
+w/o Prompts 22.48 0.696 57.64 1315.59 21.67 0.683 81.49 1542.24
+w/ Mask Codes 23.35 0.711 44.03 783.46 22.84 0.702 56.72 1067.16
+CoPeDiT 24.34 0.732 25.84 569.22 23.92 0.721 32.53 718.54
 Table 17: Quantitative results on the IXI dataset by incorporating our completeness perception prompt tokens into baseline methods instead of mask codes.
-IXI
 […]
-MMT [24] 22.64 0.698 53.60
-+ Prompts (ours) 23.19 (+0.55) 0.707 (+0.009) 46.13 (–7.47)
-Hyper-GAE [53] 22.12 0.682 72.62
-+ Prompts (ours) 22.46 (+0.34) 0.694 (+0.012) 59.43 (–13.19)
-M2DN [29] 23.47 0.715 42.52
-+ Prompts (ours) 23.84 (+0.37) 0.726 (+0.011) 33.79 (–8.73)
-C.3 Prompt Accuracy Sensitivity
-To assess how sensitive CoPeDiT is to the correctness of completeness prompts, we conduct a controlled prompt corruption study on the BraTS dataset under the number of missing modalities of 1.
-Starting from a trained CoPeDiT and CoPeVAE, we consider the degree prompt $\mathbf{p}^{d}$ (number/length of missing content), the position prompt $\mathbf{p}^{p}$ (location of missing regions), and the semantic prompt $\mathbf{p}^{s}$.
-For each prompt type, we define a corruption rate $r\in\{0,0.25,0.5,0.75,1.0\}$.
-Given a rate $r$, we randomly select an $r$-fraction of validation samples and replace the corresponding prompt with an incorrect one (sampled from other cases), while keeping all network weights and the other prompts unchanged.
-We also include an ’All’ setting where all three prompts are corrupted simultaneously.
-As illustrated in Fig.
-10, increasing the corruption rate consistently degrades synthesis quality in terms of all three metrics.
-This demonstrates that CoPeDiT relies on the accuracy of completeness prompts.
-Among the three types, corrupting the position prompt $\mathbf{p}^{p}$ causes the largest synthesis performance drop, emphasizing that precise localization of missing regions is particularly critical.
-Corrupting the degree prompt $\mathbf{p}^{d}$ also noticeably harms performance, whereas the semantic prompt $\mathbf{p}^{s}$ is somewhat more robust but still shows a clear degradation trend.
-When all prompts are corrupted, performance deteriorates the most across all metrics, confirming that the three prompts provide complementary and influential guidance for high-fidelity synthesis.
-C.4 Ablation Study
+MMT Liu et al. (2023) 22.64 0.698 53.60 1329.25
++ Prompts (ours) 23.19 (+0.55) 0.707 (+0.009) 46.13 (–7.47) 1096.06 (–233.19)
+Hyper-GAE Yang et al. (2023) 22.12 0.682 72.62 1520.49
++ Prompts (ours) 22.46 (+0.34) 0.694 (+0.012) 59.43 (–13.19) 1261.53 (–258.96)
+M2DN Meng et al. (2024) 23.47 0.715 42.52 845.29
++ Prompts (ours) 23.84 (+0.37) 0.726 (+0.011) 33.79 (–8.73) 684.62 (–160.67)
+C.2 Ablation Study
 We further evaluate the effectiveness of our prompt token design on the IXI dataset.